### v1

# Blessing of dimension in Bayesian inference on covariance matrices

Abstract
Bayesian factor analysis is routinely used for dimensionality reduction in modeling of high-dimensional covariance matrices. Factor analytic decompositions express the covariance as a sum of a low rank and diagonal matrix. In practice, Gibbs sampling algorithms are typically used for posterior computation, alternating between updating the latent factors, loadings, and residual variances. In this article, we exploit a blessing of dimensionality to develop a provably accurate pseudo-posterior for the covariance matrix that bypasses the need for Gibbs or other variants of Markov chain Monte Carlo sampling. Our proposed Factor Analysis with BLEssing of dimensionality (FABLE) approach relies on a first-stage singular value decomposition (SVD) to estimate the latent factors, and then defines a jointly conjugate prior for the loadings and residual variances. The accuracy of the resulting pseudo-posterior for the covariance improves with increasing dimensionality. We show that FABLE has excellent performance in high-dimensional covariance matrix estimation, including producing well calibrated credible intervals, both theoretically and through simulation experiments. We also demonstrate the strength of our approach in terms of accurate inference and computational efficiency by applying it to a gene expression data set.
Keywords: Bayes, Factor analysis, High-dimensional, Large p, Posterior approximation, Scalable, Singular value decomposition
1 Introduction
Inference on covariance matrices in high-dimensional data is a key focus in many application areas, motivating a rich literature on associated statistical methods. One thread of this literature avoids modeling of the data and instead focuses on high-dimensional covariance matrix estimators under various assumptions on the inherent low-dimensional structure in the data, including (but not limited to) banded covariance (Bickel and Levina,, 2008), low rank structure (Shikhaliev et al.,, 2019), low rank with sparsity (Richard et al.,, 2012), sparse covariance (Bien and Tibshirani,, 2011), and sparse inverse precision matrix estimation (Zhang and Zou,, 2014). Our interest is instead in model-based Bayesian approaches, which have advantages in terms of naturally accommodating complexities in the data and quantifying uncertainty, while having disadvantages in terms of computational efficiency, particularly in implementing Markov chain Monte Carlo (MCMC) algorithms to sample from the posterior distribution.
Our particular focus is on factor models, which express the covariance matrix in a low rank plus diagonal form, and represent a popular probabilistic model-based alternative to principal components analysis (PCA). Factor analysis introduces a lower-dimensional latent factor $\eta_{i}\in\mathbb{R}^{k}$ corresponding to each observation $y_{i}\in\mathbb{R}^{p}$ with $k\ll p$, such that $y_{i}=\Lambda\eta_{i}+\epsilon_{i}$ for a factor loadings matrix $\Lambda\in\mathbb{R}^{p\times k}$ and independent idiosyncratic errors $\epsilon_{i}\sim N_{p}(0,\Sigma)$. Following typical practice, we let the latent factors have independent Gaussian priors $\eta_{i}\sim N_{k}(0,\mathbb{I}_{k})$. Marginalizing out the latent factors leads to $y_{i}\sim N_{p}(0,\Psi)$, with the covariance matrix $\Psi=\Lambda\Lambda^{\top}+\Sigma$. This provides a convenient decomposition of the covariance, which dramatically reduces the number of free parameters required to model the $p\times p$ covariance matrix from $p(p+1)/2$ to $(p-(k-1)/2)k+p$.
There is a vibrant recent literature improving upon and expanding the scope of Bayesian factor analysis (Schiavon et al.,, 2022; De Vito et al.,, 2021; Frühwirth-Schnatter,, 2023; Roy et al.,, 2021; Ma and Liu,, 2022; Bolfarine et al.,, 2022; Xie et al.,, 2022). Even with increasingly rich classes of priors and data types, the canonical approach for posterior computation remains Gibbs samplers that iterate between updating latent factors, factor loadings, residual variances, hyperparameters controlling the hierarchical prior, and other model parameters. This approach is simple to implement in broad model classes, but commonly faces problems with slow mixing, particularly as data dimensionality and complexity increase. Issues with slow mixing and bottlenecks with large datasets have motivated alternative posterior sampling algorithms that apply Hamiltonian Monte Carlo (HMC) after marginalizing out the latent factors, exploiting sufficient statistics and parallel computation to facilitate implementation for large sample sizes (Chandra et al.,, 2023; Man and Culpepper,, 2022). Nonetheless, all current MCMC sampling algorithms for large covariance matrices face substantial computational hurdles as the number of dimensions increases.
The lack of sufficiently scalable MCMC algorithms has motivated a rich literature on developing more computationally efficient algorithms for Bayesian inference in factor models. Variational Bayes approaches (Hansen et al.,, 2023; Wang and Stephens,, 2021) attempt to approximate the posterior but often the accuracy of the approximation is poor, with a tendency to dramatically underestimate posterior uncertainty. A concise review comparing variational approximations and posterior sampling approaches for factor analysis is provided in Foo and Shim, (2021). Maximum a posteriori (MAP) estimation under sparsity priors (Srivastava et al.,, 2017) and expectation-maximization (EM) algorithms (Ročková and George,, 2016; Avalos-Pacheco et al.,, 2022; Zhao et al.,, 2016) provide computationally efficient point estimates of the covariance matrix. Sabnis et al., (2016) develop a distributed computing framework aggregating local estimates of the covariance across different cores. Srivastava et al., (2017); Ročková and George, (2016), and Sabnis et al., (2016) provide theoretical results regarding estimation accuracy and posterior concentration, without providing any guarantees on uncertainty quantification. A key open problem in the literature remains how to obtain an efficient posterior approximation for a high-dimensional covariance that has guarantees in terms of valid frequentist coverage of credible intervals.
Fast algorithms for Bayesian factor analysis capable of scaling efficiently to high-dimensional data typically sacrifice the ability to provide an accurate characterization of uncertainty. The focus of this article is on proposing a simple approach for overcoming this limitation, providing computationally efficient inference on a high-dimensional covariance matrix. With the increasing dimension, we have more and more variables loading on the same small number of latent factors. As a consequence, we obtain a blessing of dimensionality phenomenon allowing us to first pre-estimate the latent factors and then leverage on this point estimate to obtain a pseudo-posterior distribution of the factor loadings $\Lambda$ and error variances $\Sigma$ under conjugate priors on these parameters. This in turn induces a pseudo-posterior distribution for the covariance matrix $\Psi=\Lambda\Lambda^{\top}+\Sigma$. Our proposed Factor Analysis with BLEssing of dimensionality (FABLE) approach completely bypasses MCMC. To ensure that the pseudo-posterior credible intervals for the entries of the covariance are asymptotically normal with guarantees on valid frequentist coverage, we develop a data-adaptive coverage correction (CC); we refer to the adjusted FABLE approach as CC-FABLE.
Fan et al., (2023) also pre-estimate latent factors, but in the context of regression with high-dimensional correlated covariates. In their approach, a first-stage principal components analysis is carried out for the high-dimensional covariates. The results are then used to obtain a modified covariate vector consisting of principal component scores and de-correlated versions of the original covariates. Using this new covariate vector in the second stage high-dimensional regression avoids multicollinearity. This focus, and the corresponding methodology, are substantially different from ours. A blessing of dimensionality phenomenon is also explored in Li et al., (2018), where estimation of the covariance between a subset of the variables of interest is improved by incorporating auxiliary information from other variables.
In Section 2, we develop our FABLE methodology with the CC-FABLE modification and provide an approach to choose key hyperparameter values. In Section 3, we provide theoretical results on pseudo-posterior contraction rates and uncertainty quantification, leveraging on a blessing of dimensionality phenomenon. In Section 4, we validate our approach with comparison to other state-of-the-art approaches from the viewpoint of estimation error, frequentist coverage, and computational runtime in different numerical experiments. In Section 5, we present a relevant application of our approach to a large gene expression data set and highlight distinct advantages over competitors. In Section 6, we provide a discussion on future research directions.
2 Proposed Methodology
2.1 Initial Approach
The observed data consist of $\mathbf{Y}=[y_{1},\ldots,y_{n}]^{\top}\in\mathbb{R}^{n\times p}$, where $y_{i}\in\mathbb{R}^{p}$ for $i=1,\ldots,n$. We consider a factor model for $y_{i}$ given by
$y_{i}=\Lambda\eta_{i}+\epsilon_{i},\qquad\epsilon_{i}\overset{iid}{\sim}N_{p}(0,\Sigma),$ (1)
where we have omitted the intercept term, assuming the data have been centered prior to analysis. Here, $\Lambda\in\mathbb{R}^{p\times k}$ is an unknown matrix of factor loadings, $\eta_{i}\overset{iid}{\sim}N_{k}(0,\mathbb{I}_{k})$ is the $i$th latent factor, and $\epsilon_{i}$ is a zero-mean idiosyncratic error having diagonal covariance $\Sigma=\mbox{diag}(\sigma_{1}^{2},\ldots,\sigma_{p}^{2})$ for $i=1,\ldots,n$. It is typical to assume that the dimension of the latent factors $k$ is much smaller than the dimension of the data $p$. Integrating out the latent factors leads to
$y_{i}\mid\Lambda,\Sigma\overset{iid}{\sim}N_{p}(0,\Lambda\Lambda^{\top}+\Sigma).$ (2)
The covariance matrix $\Psi=\Lambda\Lambda^{\top}+\Sigma$ of the $y_{i}$s in (2) is decomposed as the sum of a low-rank and a diagonal matrix. In this paper, our goal is to estimate $L=\Lambda\Lambda^{\top},\Sigma,$ and $\Psi=\Lambda\Lambda^{\top}+\Sigma$. We first illustrate our methodology assuming $k$ is known. We discuss an approach to estimate $k$ in Section 2.3.
Let $\mathbf{M}=[\eta_{1},\ldots,\eta_{n}]^{\top}\in\mathbb{R}^{n\times k}$ and $\Lambda=[\lambda_{1},\ldots,\lambda_{p}]^{\top}\in\mathbb{R}^{p\times k}$, with $\lambda_{j}^{\top}$ denoting the $j$th row of $\Lambda$. We also denote the $j$th column of $\mathbf{Y}$ as $y^{(j)}$, so that $\mathbf{Y}=[y^{(1)},\ldots,y^{(p)}].$ The latent factor model (1) can be alternatively expressed as
$y^{(j)}=\mathbf{M}\lambda_{j}+\epsilon^{(j)},$ (3)
where $\epsilon^{(j)}$ is the $j$th column of the matrix $E=[\epsilon_{1},\ldots,\epsilon_{n}]^{\top}=[\epsilon^{(1)},\ldots,\epsilon^{(p)}],$ with $\epsilon^{(j)}\overset{ind}{\sim}N_{n}(0_{n},\sigma_{j}^{2}\mathbb{I}_{n})$ for $j=1,\ldots,p$. Writing in matrix form, we obtain
$\mathbf{Y}=\mathbf{M}\Lambda^{\top}+E.$ (4)
Our strategy to obtain computationally efficient pseudo-posterior samples for $\Psi$ is outlined as follows. We first estimate $\mathbf{M}$ by a plug-in estimator $\widehat{\mathbf{M}}$ based on the singular value decomposition (SVD) of $\mathbf{Y}$ and then consider a surrogate model to (3) where $\mathbf{M}$ is replaced by $\widehat{\mathbf{M}}$. Conditional on the observed data, $\widehat{\mathbf{M}}$ is fixed. Thus, the surrogate model is equivalent to $p$ parallel regression problems. We then let the regression coefficient and the error variance for the $j$th regression have conjugate normal-inverse gamma prior distributions for $j=1,\ldots,p$ to obtain their pseudo-posterior samples across the $p$ different regressions in parallel. Finally, we combine the pseudo-posterior samples of the regression coefficients and the error variances to obtain a pseudo-posterior sample of $\Psi$.
To motivate our choice for $\widehat{\mathbf{M}}$, we start with the SVD of $\mathbf{Y}$, given by
$\mathbf{Y}=UDV^{\top}+U_{\perp}D_{\perp}V_{\perp}^{\top},$ (5)
where $U\in\mathbb{R}^{n\times k},D\in\mathbb{R}^{k\times k},V\in\mathbb{R}^{p\times k},U_{\perp}\in\mathbb{R}^{n\times(r-k)},D_{\perp}\in\mathbb{R}^{(r-k)\times(r-k)},$ and $V_{\perp}\in\mathbb{R}^{p\times(r-k)}$, with $r=p\wedge n.$ The columns of $U,U_{\perp},V,V_{\perp}$ consist of orthonormal vectors, with $U^{\top}U_{\perp}=V^{\top}V_{\perp}=\mathbf{O}_{k\times(r-k)}$, where $\mathbf{O}_{p_{1}\times p_{2}}$ denotes the $p_{1}\times p_{2}$ matrix with all entries equal to 00. Let $\mathbf{A}=\mathbf{Y}V/\sqrt{p}=UD/{\sqrt{p}}$ and $\widehat{C}\in\mathbb{R}^{k\times k}$ be an invertible matrix satisfying
$\widehat{C}\widehat{C}^{\top}=\dfrac{1}{n}\mathbf{A}^{\top}\mathbf{A}=\dfrac{D^{2}}{np}.$ (6)
An immediate choice of $\widehat{C}$ satisfying (6) is given by $\widehat{C}=D/\sqrt{np}$. Lastly, we let
$\widehat{\mathbf{M}}=\mathbf{A}(\widehat{C}^{\top})^{-1}.$ (7)
For the choice of $\widehat{C}=D/\sqrt{np},$ it is straightforward from (7) that $\widehat{\mathbf{M}}=\sqrt{n}U.$ We will denote this particular choice of $\widehat{\mathbf{M}}=\sqrt{n}U$ as the canonical choice of $\widehat{\mathbf{M}}$, observing that other choices are also possible for different choices of $\widehat{C}$ satisfying (6). We provide a heuristic leading to (7) at the end of this Section.
Given a particular choice of $\widehat{\mathbf{M}}$, we consider the following surrogate model:
$y^{(j)}=\widehat{\mathbf{M}}\widetilde{\lambda}_{j}+\widetilde{\epsilon}^{(j)},\quad\widetilde{\epsilon}^{(j)}\overset{ind}{\sim}N_{n}(0,\widetilde{\sigma}_{j}^{2}\mathbb{I}_{n})$ (8)
for $j=1,\ldots,p$. The model (8) could be interpreted as a version of (3) with $\widehat{\mathbf{M}}$ substituted for the original matrix of latent factors $\mathbf{M}$ and new parameters $\widetilde{\lambda}_{j}\in\mathbb{R}^{k}$, $\widetilde{\sigma}_{j}^{2}>0$ for $j=1,\ldots,p$. Next, we endow $(\widetilde{\lambda}_{j},\widetilde{\sigma}_{j}^{2})_{j=1}^{p}$ with independent normal-inverse gamma (NIG) priors $(\widetilde{\lambda}_{j},\widetilde{\sigma}_{j}^{2})\overset{iid}{\sim}\mbox{NIG}(0_{k},\tau^{2}\mathbb{I}_{k},\gamma_{0}/2,\gamma_{0}\delta_{0}^{2}/2)$. That is, we let
$\widetilde{\lambda}_{j}\mid\widetilde{\sigma}_{j}^{2}$ $\sim N_{k}(0_{k},\widetilde{\sigma}_{j}^{2}\,\tau^{2}\,\mathbb{I}_{k}),\quad\widetilde{\sigma}_{j}^{2}\sim\mbox{IG}\left(\dfrac{\gamma_{0}}{2},\dfrac{\gamma_{0}\delta_{0}^{2}}{2}\right).$ (9)
The global shrinkage parameter $\tau^{2}$ allows us to a priori shrink the factor loadings towards zero, regularizing $\widetilde{\Lambda}=[\widetilde{\lambda}_{1},\ldots,\widetilde{\lambda}_{p}]^{\top}$ and favoring sparsity. We discuss a data-driven strategy to estimate $\tau^{2}$ in Section 2.3.
The surrogate model (8) and the conjugate prior specification (9) motivate the pseudo-posterior densities $\widetilde{\Pi}_{j}$ of $(\widetilde{\lambda}_{j},\widetilde{\sigma}_{j}^{2})$ for each $j=1,\ldots,p$, given by
$\begin{split}\widetilde{\Pi}_{j}(\widetilde{\lambda}_{j},\widetilde{\sigma}_{j}^{2})&=\mbox{NIG}\left(\widetilde{\lambda}_{j},\widetilde{\sigma}_{j}^{2}\mid\mu_{j},\mathbf{K},\gamma_{n}/2,\gamma_{n}\delta_{j}^{2}/2\right)\\ &=N_{k}\left(\widetilde{\lambda}_{j}\mid\mu_{j},\widetilde{\sigma}_{j}^{2}\mathbf{K}\right)\mbox{IG}\left(\widetilde{\sigma}_{j}^{2}\,\middle|\,\dfrac{\gamma_{n}}{2},\dfrac{\gamma_{n}\delta_{j}^{2}}{2}\right),\end{split}$ (10)
with the updated hyperparameters given by
$\begin{split}\mu_{j}&=\left(\widehat{\mathbf{M}}^{\top}\widehat{\mathbf{M}}+\dfrac{\mathbb{I}_{k}}{\tau^{2}}\right)^{-1}\widehat{\mathbf{M}}^{\top}y^{(j)},\\ \mathbf{K}&=\left(\widehat{\mathbf{M}}^{\top}\widehat{\mathbf{M}}+\dfrac{\mathbb{I}_{k}}{\tau^{2}}\right)^{-1},\\ \gamma_{n}&=\gamma_{0}+n,\\ \gamma_{n}\delta_{j}^{2}&=\gamma_{0}\delta_{0}^{2}+\left(y^{(j)\top}y^{(j)}-\mu_{j}^{\top}\mathbf{K}^{-1}\mu_{j}\right).\end{split}$ (11)
The pseudo-posteriors are simply obtained by treating $\widehat{\mathbf{M}}$ as a fixed design matrix for the $j$th surrogate regression and then applying Bayes’ rule. To obtain independent pseudo-posterior samples of $\Psi$, we first draw independent samples $(\widetilde{\lambda}_{j},\widetilde{\sigma}_{j}^{2})\overset{ind}{\sim}\widetilde{\Pi}_{j}$ for $j=1,\ldots,p$. We then let $\widetilde{\Lambda}=[\widetilde{\lambda}_{1},\ldots,\widetilde{\lambda}_{p}]^{\top},\widetilde{L}=\widetilde{\Lambda}\widetilde{\Lambda}^{\top},\widetilde{\Sigma}=\mbox{diag}(\widetilde{\sigma}_{1}^{2},\ldots,\widetilde{\sigma}_{p}^{2}),$ and $\widetilde{\Psi}=\widetilde{L}+\widetilde{\Sigma}$ denote the pseudo-posterior samples of $\Lambda,L=\Lambda\Lambda^{\top}$, $\Sigma$, and $\Psi$, respectively. Our proposed Factor Analysis with BLEssing of dimensionality (FABLE) approach obtains independent pseudo-posterior samples of the covariance matrix $\Psi$ in an embarrassingly parallel fashion, bypassing the need to carry out MCMC.
A natural concern is whether different choices of $\widehat{\mathbf{M}}$ satisfying (7) would affect the distribution of $\widetilde{\Psi}$. The following result, with proof in Section A of the Supplementary Material, ensures that this is not the case.
Proposition 2.1.
(i) For any $\widehat{\mathbf{M}}$ satisfying (7), we have $\widehat{\mathbf{M}}^{\top}\widehat{\mathbf{M}}=n\mathbb{I}_{k}$ and $\widehat{\mathbf{M}}\widehat{\mathbf{M}}^{\top}=nUU^{\top}.$ (ii) The pseudo-posterior distribution of $\widetilde{L},\widetilde{\Sigma},$ and $\widetilde{\Psi}$ obtained from the FABLE approach only depends on $\widehat{\mathbf{M}}$ through $\widehat{\mathbf{M}}^{\top}\widehat{\mathbf{M}}$ and $\widehat{\mathbf{M}}\widehat{\mathbf{M}}^{\top}$.
In our experience, the pseudo-posterior samples $\widetilde{\Psi}$ perform well when estimating the true covariance matrix, both in terms of simulation studies and theoretical posterior contraction rates. However, we observed that the entry-wise pseudo-posterior credible intervals obtained from the samples $\widetilde{\Psi}$ tend to underestimate the uncertainty associated with estimating the entry-wise elements of the true covariance matrix, from the perspective of frequentist coverage. In Section 2.2, we build on the initial FABLE approach developed in this subsection and propose a coverage correction (CC) of FABLE called CC-FABLE. We provide provably accurate guarantees on posterior contraction rates and frequentist coverage of the CC-FABLE procedure in Section 3.
Lastly, we describe the heuristic leading to (7). From (4), the matrix $\mathbf{A}=\mathbf{Y}V/\sqrt{p}$ satisfies
$\mathbf{A}=\mathbf{M}\dfrac{\Lambda^{\top}V}{\sqrt{p}}+\dfrac{EV}{\sqrt{p}}.$
Let $C=V^{\top}\Lambda/\sqrt{p}\in\mathbb{R}^{k\times k}$. Based on the consistency of spectral estimates, we expect $E$ to be approximately independent of $V$ as both $n,p$ grow. As a result, we expect $EV/\sqrt{p}\approx 0$ for increasing $p$. This leads us to
$\mathbf{A}\approx\mathbf{M}C^{\top}$ (12)
or equivalently, $a_{i}\approx C\eta_{i}$ for $i=1,\ldots,n$, where $a_{i}$ is the $i$th row of $\mathbf{A}$ for $i=1,\ldots,n$. Since $\eta_{i}\sim N_{k}(0,\mathbb{I}_{k}),$ the marginal density of $a_{i}$ is approximately given by $a_{i}\overset{iid}{\sim}N_{k}(0_{k},CC^{\top})$, for $i=1,\ldots,n$. This motivates the following estimator $\widehat{C}\widehat{C}^{\top}$ of $CC^{\top}$:
$\widehat{C}\widehat{C}^{\top}=\dfrac{1}{n}\mathbf{A}^{\top}\mathbf{A}=\dfrac{1}{np}D^{2},$
leading us to (6). Given any $\widehat{C}$ satisfying (6), we use (12) to propose $\widehat{\mathbf{M}}=\mathbf{A}(\widehat{C}^{\top})^{-1}$ as an estimate of $\mathbf{M}.$ In particular, the canonical choice $\widehat{\mathbf{M}}=\sqrt{n}U$ is the same as the estimator obtained by carrying out principal components analysis (PCA) on the matrix $\mathbf{Y}\mathbf{Y}^{\top}$ (Fan et al.,, 2023). However, (7) permits other choices corresponding to different choices of $\widehat{C}$ satisfying (6), providing a general framework for obtaining estimates of the latent factors.
2.2 Coverage Correction
Although the pseudo-posterior of $\widetilde{\Psi}$ obtained from (10) performed well in terms of point estimation across simulation studies, we observed under-coverage of the entry-wise credible intervals, particularly when the signal-to-noise ratios $R_{j}=\|\lambda_{j}\|_{2}^{2}/\sigma_{j}^{2}$ for $j=1,\ldots,p$ are large. To improve coverage, we propose a simple coverage correction (CC) of the FABLE approach, which we denote by CC-FABLE. We provide a theoretical justification of this approach in Section 3. We found that the correction leads to improved performance in simulation studies, particularly with increasing $R_{j}$s.
To improve coverage, we include an additional factor $\rho^{2}$ in the conditional pseudo-posterior variance when generating $\widetilde{\lambda}_{j}$ for $j=1,\ldots,p$ following (10). That is, we modify $\widetilde{\Pi}_{j}$ in (10) to
${\Pi}_{j}^{C}\left(\lambda_{C,j},\sigma_{C,j}^{2}\right)=N_{k}\left(\lambda_{C,j}\middle|\mu_{j},\rho^{2}\sigma_{C,j}^{2}\mathbf{K}\right)\mbox{IG}\left(\sigma_{C,j}^{2}\middle|\dfrac{\gamma_{n}}{2},\dfrac{\gamma_{n}\delta_{j}^{2}}{2}\right),$ (13)
for $j=1,\ldots,p$, where $\mathbf{K}=\{\mathbf{M}^{\top}\mathbf{M}+(\mathbb{I}_{k}/\tau^{2})\}^{-1}=\mathbb{I}_{k}/(n+\tau^{-2})$ and $\gamma_{n},\mu_{j},\delta_{j}^{2}$ are as in (11). We then proceed as before to draw coverage-corrected pseudo-posterior samples $(\lambda_{C,j},\sigma_{C,j}^{2})\overset{ind}{\sim}\Pi_{j}^{C}$ for $j=1,\ldots,p$. Letting $\Lambda_{C}=[\lambda_{C,1},\ldots,\lambda_{C,p}]^{\top}$ and $\Sigma_{C}=\mbox{diag}(\sigma_{C,1}^{2},\ldots,\sigma_{C,p}^{2})$, we construct a coverage-corrected (CC) pseudo-posterior sample of the covariance matrix as $\Psi_{C}=\Lambda_{C}\Lambda_{C}^{\top}+\Sigma_{C}.$ Having computed the hyperparameters $\gamma_{n}$ and $\delta_{j}^{2}$ for $j=1,\ldots,p$ as in (11), the pseudo-posterior mean $\widehat{\Psi}_{C}$ of $\Psi_{C}$ is available analytically, with its $uv$th entry for $1\leq u,v\leq p$ given by
$\begin{split}\widehat{\Psi}_{C,uv}&=\left\{\begin{array}[]{ll}\mu_{u}^{\top}\mu_{v},&\text{if $u\neq v,$}\\ \|\mu_{u}\|_{2}^{2}+\left(1+\dfrac{k\rho^{2}}{n+\tau^{-2}}\right)\dfrac{\gamma_{n}\delta_{u}^{2}}{\gamma_{n}-2},&\text{if $u=v,$}\end{array}\right.\end{split}$ (14)
provided $\gamma_{n}>2$. For finite $k,\rho^{2},$ and $\tau^{2}$, the pseudo-posterior mean $\widehat{\Psi}_{C}$ can approximated as $\widehat{\Psi}_{C}=G_{0}G_{0}^{\top}+\Delta,$ where $G_{0}=[\mu_{1},\ldots,\mu_{p}]^{\top}$ and $\Delta=\mbox{diag}(\delta_{1}^{2},\ldots,\delta_{p}^{2})$, as $n$ increases. From here on, we take $\widehat{\Psi}_{C}=G_{0}G_{0}^{\top}+\Delta$ to be the pseudo-posterior mean of $\Psi$ under CC-FABLE, unless mentioned otherwise.
We choose $\rho$ to ensure that a summary statistic of the entrywise asymptotic coverages is $1-\alpha$, such as their average or minimum. Let $B$ denote the $p\times p$ symmetric matrix with entries $b_{uv}$, where
$\begin{split}b_{uv}&=\left\{\begin{array}[]{ll}\left(1+\dfrac{\|\mu_{u}\|_{2}^{2}\|\mu_{v}\|_{2}^{2}+(\mu_{u}^{\top}\mu_{v})^{2}}{\mathcal{V}_{u}^{2}\|\mu_{v}\|^{2}+\mathcal{V}_{v}^{2}\|\mu_{u}\|_{2}^{2}}\right)^{1/2},&\text{if $u\neq v,$}\\ \left(1+\dfrac{\|\mu_{u}\|_{2}^{2}}{2\mathcal{V}_{u}^{2}}\right)^{1/2},&\text{if $u=v,$}\end{array}\right.\end{split}$ (15)
with $\mathcal{V}_{j}^{2}=\|(\mathbb{I}_{n}-UU^{\top})y^{(j)}\|_{2}^{2}/n$ for $j=1,\ldots,p$. For example, we can ensure that all the entrywise asymptotic coverages are at least $1-\alpha$ by letting $\rho=\sup_{u,v}b_{uv}$. Another possibility is to let $\rho$ be the root of a nonlinear equation, which ensures the average of the asymptotic entrywise coverages is $1-\alpha$; we discuss the associated nonlinear equation along with the rationale behind possible choices of $\rho$ at the end of Section 3.3. Based on our experiences with finite sample simulations, we recommend choosing $\rho$ to the mean of $b_{uv}$ for $1\leq u\leq v\leq p$ as a computationally efficient strategy. This performed well with both median and average entrywise coverages close to $1-\alpha$. In what follows, we only consider the CC-FABLE version of our approach for theoretical justification, simulation results, and the real data application.
2.3 Hyperparameter Choice
Before obtaining pseudo-posterior samples of the covariance matrix, one needs to carefully choose two key hyperparameters. These are the number of factors $k$ and the global variance of the factor loadings $\tau^{2}$. To estimate $k$, we use the approach described in Chen and Li, (2022). The authors construct an information criterion $\mbox{JIC}(k)$ by adding the penalty $\nu(k,n,p)=k(n\vee p)\log(n\wedge p)$ to the model deviance. The proposed $\mbox{JIC}(k)$ criterion is similar in spirit to the Akaike information criterion (AIC) or the Bayesian information criterion (BIC) (Akaike,, 1974; Schwarz,, 1978). Given an upper bound $\mathcal{K}_{0}$ for the possible values of $k$, we estimate $\widehat{k}=\min_{1\leq k\leq\mathcal{K}_{0}}\mbox{JIC}(k).$ To choose an upper bound $\mathcal{K}_{0}$ for plausible values of $k$, we consider the singular values $s_{1}\geq s_{2}\geq\ldots\geq s_{n\wedge p}$ of $\mathbf{Y}$ and let $\mathcal{K}_{0}$ be the smallest integer $\mathcal{K}$ such that $\left(\sum_{j=1}^{\mathcal{K}}s_{j}\right)/\left(\sum_{j=1}^{n\wedge p}s_{j}\right)\geq S_{0}$, where we typically let $S_{0}$ be $0.5$ or $0.75$. In our experience across simulation studies and applications to gene expression datasets, we found this choice of $\widehat{k}$ to perform well.
We next proceed to estimate $\tau^{2}$, which can be interpreted as the global shrinkage parameter of the loadings. For this, we employ an empirical Bayes approach. Since $\widetilde{\lambda}_{j}\mid\widetilde{\sigma}_{j}^{2},\tau^{2}\sim N_{k}(0,\tau^{2}\widetilde{\sigma}_{j}^{2}\mathbb{I}_{k})$ for $j=1,\ldots,p$ a priori, we obtain an estimate of $\tau^{2}$ by conditioning on $(\widetilde{\lambda}_{j},\widetilde{\sigma}_{j}^{2})$ and then maximizing the conditional likelihood, leading to $\widehat{\tau}_{0}^{2}=\sum_{j=1}^{p}\left({\|\widetilde{\lambda}_{j}\|_{2}^{2}}/{\widetilde{\sigma}_{j}^{2}}\right)/(kp).$ Following from the pseudo-posterior updates (10) and (11), we simply estimate $k$ by $\widehat{k}$ as before, $\|\widetilde{\lambda}_{j}\|_{2}^{2}$ by $\mathcal{L}_{j}^{2}=\|U^{\top}y^{(j)}\|_{2}^{2}/n$, and $\widetilde{\sigma}_{j}^{2}$ by $\mathcal{V}_{j}^{2}=\|(\mathbb{I}_{n}-UU^{\top})y^{(j)}\|_{2}^{2}/n$ for $j=1,\ldots,p$. We remark here that $U$ is dependent on the estimate $\widehat{k}$. This leads us to the plug-in estimate of $\tau^{2}$ given by
$\widehat{\tau}^{2}=\dfrac{1}{p\widehat{k}}\sum_{j=1}^{p}\dfrac{\mathcal{L}_{j}^{2}}{\mathcal{V}_{j}^{2}}.$ (16)
The consistency of the point estimators $\mathcal{L}_{j}^{2}$ and $\mathcal{V}_{j}^{2}$ for $1\leq j\leq p$ is proven in Section E, Lemma E.4 of the Supplementary Material. Lastly, we observed that inference using the CC-FABLE procedure is not sensitive to $(\gamma_{0},\delta_{0}^{2})$, and thus we use the default values $\gamma_{0}=\delta_{0}^{2}=1$ in practice. Once the number of factors $k$ and the common variance $\tau^{2}$ have been estimated using $\widehat{k}$ and $\widehat{\tau}^{2}$, respectively, we then proceed to obtain pseudo-posterior samples of $\Psi$ using CC-FABLE as described before.
2.4 Final Algorithm
In this subsection, we provide an algorithm for implementing CC-FABLE on a given data set to obtain pseudo-posterior samples of the high-dimensional covariance matrix. We first tune the relevant hyperparameters $k$ and $\tau^{2}$ as in Section 2.3 and proceed to obtain coverage-corrected samples as in Section 2.2. The relevant steps are described in Algorithm 1.
Input: The data matrix $\mathbf{Y}\in\mathbb{R}^{n\times p}$, number of Monte Carlo (MC) samples $N_{0}$, inverse-gamma hyperparameters $(\gamma_{0},\sigma_{0}^{2}),$ and the upper bound on the cumulative singular value proportion ${S}_{0}$. Let $r=n\wedge p$.
Step 1: Compute the SVD of $\mathbf{Y}$ as $\mathbf{Y}=U^{*}D^{*}V^{*\top}$ with $U\in\mathbb{R}^{n\times r},D\in\mathbb{R}^{r\times r},V\in\mathbb{R}^{p\times r}$, and $D=\mbox{diag}(s_{1},\ldots,s_{r})$ such that $s_{1}\geq s_{2}\geq\ldots\geq s_{r}\geq 0.$ Step 2: Let $\widehat{k}={\mbox{argmin}}\,\{\mbox{JIC}(k^{*})\,|\,k^{*}=1,\ldots,\mathcal{K}_{0}\},$ where $\mathcal{K}_{0}=\mbox{argmin}_{1\leq\mathcal{K}\leq r}\left(\sum_{j=1}^{\mathcal{K}}s_{j}\right)/\left(\sum_{j=1}^{r}s_{j}\right).$ Step 3: Let $U$ consist of the first $\widehat{k}$ columns of $U^{*}$. For $1\leq j\leq p$, let $y^{(j)}$ denote the $j$th column of $\mathbf{Y}$, and obtain $\mathcal{L}_{j}^{2}=\|U^{\top}y^{(j)}\|_{2}^{2}/n$ and $\mathcal{V}_{j}^{2}=\|(\mathbb{I}_{n}-UU^{\top})y^{(j)}\|_{2}^{2}/n.$ Step 4: Estimate $\tau^{2}$ by
$\widehat{\tau}^{2}=\dfrac{1}{p\widehat{k}}\sum_{j=1}^{p}\dfrac{\mathcal{L}_{j}^{2}}{\mathcal{V}_{j}^{2}}.$
Step 5: For $1\leq j\leq p$, let $\mu_{j}=\sqrt{n}U^{\top}y^{(j)}/\left(n+\widehat{\tau}^{-2}\right).$ For $1\leq u,v\leq p$, let
$\begin{split}b_{uv}&=\left\{\begin{array}[]{ll}\left(1+\dfrac{\|\mu_{u}\|_{2}^{2}\|\mu_{v}\|_{2}^{2}+(\mu_{u}^{\top}\mu_{v})^{2}}{\mathcal{V}_{u}^{2}\|\mu_{v}\|^{2}+\mathcal{V}_{v}^{2}\|\mu_{u}\|_{2}^{2}}\right)^{1/2},&\text{if $u\neq v,$}\\ \left(1+\dfrac{\|\mu_{u}\|_{2}^{2}}{2\mathcal{V}_{u}^{2}}\right)^{1/2},&\text{if $u=v,$}\end{array}\right.\end{split}$
and let $\rho=\overline{b}=\left(\sum_{1\leq u\leq v\leq p}b_{uv}\right)/\{p(p+1)/2\}$ be the coverage-correction factor. Step 6: Let $\gamma_{n}=\gamma_{0}+n$. For $j=1,\ldots,p$, evaluate
$\gamma_{n}\delta_{j}^{2}=\gamma_{0}\delta_{0}^{2}+y^{(j)\top}\left(\mathbb{I}_{n}-\dfrac{nUU^{\top}}{n+\widehat{\tau}^{-2}}\right)y^{(j)}.$
Step 7: For each $t=1,\ldots,N_{0}$, independently sample $({\lambda}_{C,j}^{(t)},{\sigma}_{C,j}^{(t)2})$ for $j=1,\ldots,p$ from
${\sigma}_{C,j}^{(t)2}\sim\mbox{IG}\left(\dfrac{\gamma_{n}}{2},\dfrac{\gamma_{n}\delta_{j}^{2}}{2}\right),\qquad{\lambda}_{C,j}^{(t)}\mid{\sigma}_{C,j}^{(t)2}\sim N_{k}\left(\mu_{j},\dfrac{\rho^{2}{\sigma}_{C,j}^{(t)2}}{n+\widehat{\tau}^{-2}}\mathbb{I}_{k}\right).$
Form ${\Lambda}_{C}^{(t)}=[{\lambda}_{C,1}^{(t)},\ldots,{\lambda}_{C,p}^{(t)}]^{\top}$ and ${\Sigma}_{C}^{(t)}=\mbox{diag}({\sigma}_{C,1}^{(t)2},\ldots,{\sigma}_{C,p}^{(t)2})$. Step 8: For each $t=1,\ldots,N_{0}$, let $L_{C}^{(t)}=\Lambda_{C}^{(t)}\Lambda_{C}^{(t)\top}$, and compute the $t$th coverage-corrected sample as
$\Psi_{C}^{(t)}=L_{C}^{(t)}+\Sigma_{C}^{(t)}.$
Output: The $N_{0}$ MC samples of the covariance matrix $\Psi_{C}^{(1)},\ldots,\Psi_{C}^{(N_{0})}$.
Algorithm 1 Steps to obtain $N_{0}$ pseudo-posterior samples of the covariance matrix using CC-FABLE.
3 Theoretical Support
3.1 Setup and Assumptions
In this section, we provide theoretical guarantees of the proposed CC-FABLE procedure in estimating the high-dimensional covariance matrix. Most of the existing literature (Pati et al.,, 2014; Bhattacharya and Dunson,, 2011) on obtaining provable guarantees for Bayesian factor models utilizes the machinery of Bayes’ rule in deriving posterior contraction rates. There is a clear need for further theoretical work on uncertainty quantification in such models. Xie and Wu, (2023) consider a signal-plus-noise model and derive theoretical properties of a quasi-posterior based approach, with guarantees on asymptotic frequentist coverage of the quasi-posterior intervals. However, their model only considers a square signal matrix, with the underlying focus different from high-dimensional covariance estimation. In contrast, the proposed method carries out inference on the underlying covariance matrix, without directly relying on Bayes’ rule to obtain the pseudo-posterior samples, which makes theoretical analysis of the proposed approach challenging. We overcome this challenge by leveraging on a blessing of dimensionality phenomenon which requires both $n$ and $p=p_{n}$ to grow, providing results on both the pseudo-posterior contraction and uncertainty quantification of our approach. The blessing of dimensionality phenomenon allows accurate estimation of the latent factor subspace up to rotational ambiguity, which serves as an important component behind the theoretical results. The proofs of all the theorems can be found in Sections B-D of the Supplementary Material.
We assume the following data-generating model:
$y_{i}=\Lambda_{0}\eta_{0i}+\epsilon_{i},$ (17)
where $\epsilon_{i}\overset{iid}{\sim}N_{p}(0,\Sigma_{0})$ for $i=1,\ldots,n$ with $\Sigma_{0}=\mbox{diag}(\sigma_{01}^{2},\ldots,\sigma_{0p}^{2}),$ and $\eta_{0i}\overset{iid}{\sim}N_{k}(0,\mathbb{I}_{k})$ for $i=1,\ldots,n$. Here, $\Lambda_{0}=[\lambda_{01},\ldots,\lambda_{0p}]^{\top}$ is the true factor loadings matrix while $\eta_{0i}$ are the true latent factors; integrating them out provides us $y_{i}\overset{iid}{\sim}N_{p}(0,\Lambda_{0}\Lambda_{0}^{\top}+\Sigma_{0})$ for $i=1,\ldots,n$ as the marginal distribution of the data. Let $\mathbf{Y}=[y_{1},\ldots,y_{n}]^{\top}\in\mathbb{R}^{n\times p}$ and $M_{0}=[\eta_{01},\ldots,\eta_{0n}]^{\top}\in\mathbb{R}^{n\times k}$ be the data matrix and the true matrix of latent factors, respectively, so that the true data generating model may be written as
$\mathbf{Y}=M_{0}\Lambda_{0}^{\top}+E,$ (18)
where $E=[\epsilon_{1},\ldots,\epsilon_{n}]^{\top}.$ Our primary goal is the inference for the covariance matrix $\Psi_{0}=\Lambda_{0}\Lambda_{0}^{\top}+\Sigma_{0}.$
For a matrix $A\in\mathbb{R}^{n_{1}\times n_{2}}$, suppose the singular values of $A$ are given by $s_{1}(A)\geq\ldots\geq s_{n_{1}\wedge n_{2}}(A)$; let $\|A\|=\underset{\|x\|_{2}=1}{\sup}\|Ax\|_{2}=s_{1}(A)$ denote the operator norm of $A$ and $\|A\|_{\infty}=\max_{ij}|A_{ij}|$ denote the max-norm of $A$, with $A_{ij}$ denoting the $ij$th entry of $A$. For two non-negative sequences $a_{m},b_{m}$, we say $a_{m}\asymp b_{m}$ if $a_{m}=\mathcal{O}(b_{m})$ and $b_{m}=\mathcal{O}(a_{m})$ as $m\to\infty$. We assume the following conditions on the true data-generating model:
Assumption 3.1.
We have $p_{n}\to\infty$ and $(\log p_{n})/n=o(1)$ as $n\to\infty$.
Assumption 3.2.
$s_{k}(\Lambda_{0})\asymp\|\Lambda_{0}\|\asymp\sqrt{p_{n}}$ as $n\to\infty$, $\|\Lambda_{0}\|_{\infty}<\infty$, and $\underset{1\leq j\leq p_{n}}{\min}\|\lambda_{0j}\|_{2}\geq c_{1}$ for some constant $c_{1}>0$.
Assumption 3.3.
The true error variances satisfy $\underset{1\leq j\leq p_{n}}{\max}\sigma_{0j}^{2}=\mathcal{O}(1)$ and $\underset{1\leq j\leq p_{n}}{\min}\sigma_{0j}^{2}\geq c_{2}$ for some constant $c_{2}>0$.
Assumption 3.4.
The hyperparameters $k,\tau^{2},\gamma_{0},\delta_{0}^{2}$, and $\rho$ are fixed constants.
Such assumptions are standard in the existing literature on asymptotic theoretical properties of latent factor models. Assumption 3.1 allows the number of dimensions $p_{n}$ to scale as any polynomial function of $n$. Assumption 3.2 ensures that the true loadings matrix $\Lambda_{0}$ is well-conditioned, with the low-rank portion $\Lambda_{0}\Lambda_{0}^{\top}$ identifiable from noise in the asymptotic regime, and also ensures that none of the columns of $\mathbf{Y}$ consist purely of noise. Assumptions 3.3 and 3.4 assume the scalar error variances and model hyperparameters are finite. For our theoretical requirements, we will assume that $k$ and $\tau^{2}$ are known and fixed. To ease exposition, we will often suppress the dependence of $p=p_{n}$ on $n$ and express it when needed.
Let $M_{0}\Lambda_{0}^{\top}=U_{0}D_{0}V_{0}^{\top}$ be the singular value decomposition of the signal, with $U_{0}\in\mathbb{R}^{n\times k},V_{0}\in\mathbb{R}^{p\times k}$ having orthonormal columns and $D_{0}\in\mathbb{R}^{k\times k}$ a diagonal matrix of positive singular values. Suppose the singular value decomposition of $\mathbf{Y}$ is
$\mathbf{Y}=UDV^{\top}+U_{\perp}D_{\perp}V_{\perp}^{\top},$
where $U\in\mathbb{R}^{n\times k},V\in\mathbb{R}^{p\times k}$ have orthonormal columns, and $D\in\mathbb{R}^{k\times k}$ contain the $k$ largest singular values of $\mathbf{Y}$. We first provide a result showcasing the blessing of dimensionality when estimating $U_{0}U_{0}^{\top}$ by $UU^{\top}$, which forms a key part of the results that follow. Let us denote the induced CC-FABLE pseudo-posterior measure, the true data-generating measure, and the expectation under the true data-generating measure by ${\Pi}_{C}$, $P_{0}$, and $E_{0}$, respectively.
Proposition 3.5.
Suppose Assumptions 3.13.4 hold. Then, there exists a constant $G_{1}>0$ such that
$\lim_{n\to\infty}P_{0}\left\{\|UU^{\top}-U_{0}U_{0}^{\top}\|>G_{1}\left(\dfrac{1}{n}+\dfrac{1}{p_{n}}\right)\right\}=0.$
The proof of Proposition 3.5 is available in Section E.1 of the Supplementary Material. First, in Theorem 3.6, we provide psuedo-posterior contraction rates when estimating $\Lambda_{0}\Lambda_{0}^{\top},\Sigma_{0},$ and $\Psi_{0}$ using $L_{C},{\Sigma}_{C}$, and ${\Psi}_{C}=L_{C}+\Sigma_{C}$, respectively, obtained from the CC-FABLE methodology in Section 2.2. Later, using Theorems 3.7 and 3.8, we justify the uncertainty quantification of the entrywise elements of $\Psi_{0}$ via pseudo-posterior credible intervals obtained from CC-FABLE.
3.2 Pseudo-posterior Contraction Rates
Theorem 3.6.
Suppose Assumptions 3.13.4 hold. Then, as $n\to\infty$,
  1. (a)
    There exists a constant $D_{1}>0$ such that
    $E_{0}\left[{\Pi}_{C}\left\{\dfrac{\|L_{C}-\Lambda_{0}\Lambda_{0}^{\top}\|}{\|\Lambda_{0}\Lambda_{0}^{\top}\|}>D_{1}\left(\dfrac{1}{\sqrt{n}}+\dfrac{1}{\sqrt{p_{n}}}\right)\right\}\right]\to 0.$
  2. (b)
    There exists a constant $D_{2}>0$ such that
    $E_{0}\left[{\Pi}_{C}\left\{\|\Sigma_{C}-\Sigma_{0}\|>D_{2}\left(\left(\dfrac{\log p_{n}}{n}\right)^{1/3}+\dfrac{1}{p_{n}}\right)\right\}\right]\to 0.$
  3. (c)
    There exists a constant $D>0$ such that
    $E_{0}\left[{\Pi}_{C}\left\{\dfrac{\|\Psi_{C}-\Psi_{0}\|}{\|\Psi_{0}\|}>D\left(\dfrac{1}{\sqrt{n}}+\dfrac{1}{\sqrt{p_{n}}}\right)\right\}\right]\to 0.$
Theorem 3.6 shows pseudo-posterior concentration of the relevant quantities of the CC-FABLE procedure. The relative errors when estimating each of these true parameters of interest converge to 00 at the rate $n^{-1/2}+p_{n}^{-1/2}$, which showcases the blessing of dimensionality acting on the FABLE procedure. One obtains pseudo-posterior concentration if and only if both the number of samples $n$ and the number of dimensions $p_{n}$ increase. The proof of Theorem 3.6 is available in Section B of the Supplementary Material.
3.3 Uncertainty Quantification
We now consider uncertainty quantification of entrywise elements of the true covariance matrix $\Psi_{0}$. For $1\leq u,v\leq p,$ let ${\Psi}_{C,uv}$ and $\Psi_{0,uv}$ denote the $uv$th element of ${\Psi}_{C}$ and $\Psi_{0}$, respectively, where $\Psi_{C}$ denotes the coverage-corrected pseudo-posterior sample obtained from the CC-FABLE discussed in Section 2.2. For $1\leq u,v\leq p$, we have $\Psi_{0,uv}=\lambda_{0u}^{\top}\lambda_{0v}+\sigma_{0u}^{2}\mathbbm{1}(u=v)$, where $\mathbbm{1}$ denotes the indicator function. In what follows, we discuss results for a fixed value of the coverage-correction factor $\rho>0$ and then provide an approach to appropriately choose $\rho$.
We now let $T_{uv}=\mu_{u}^{\top}\mu_{v}+\delta_{u}^{2}\mathbbm{1}(u=v)$ be an estimator of $\Psi_{0,uv}$. Then, we have the following result approximating the pseudo-posterior distribution of ${\Psi}_{C,uv}$ by a suitable Gaussian distribution centered at $T_{uv}$ as both $n,p_{n}$ increase with $\rho>0$ fixed.
Theorem 3.7.
Suppose Assumptions 3.13.4 hold and let $\sqrt{n}/p_{n}=o(1)$. For $1\leq u,v\leq p$, let
$\begin{split}l_{0,uv}^{2}(\rho)&=\left\{\begin{array}[]{ll}\rho^{2}(\sigma_{0v}^{2}\|\lambda_{0u}\|_{2}^{2}+\sigma_{0u}^{2}\|\lambda_{0v}\|_{2}^{2}),&\text{for $u\neq v$},\\ 2\sigma_{0u}^{4}+4\rho^{2}\,\sigma_{0u}^{2}\|\lambda_{0u}\|_{2}^{2},&\text{for$u=v$.}\end{array}\right.\end{split}$
Then, as $n\to\infty,$
$\sup_{x\in\mathbb{R}}\,\left|{\Pi}_{C}\left\{\dfrac{\sqrt{n}({\Psi}_{C,uv}-T_{uv})}{\l_{0,uv}(\rho)}\leq x\right\}-\Phi(x)\right|\overset{P_{0}}{\to}0\quad\text{for all $1\leq u,v\leq p.$}$
Theorem 3.7 allows us to approximate the asymptotic pseudo-posterior distribution of each element of the covariance matrix after suitable centering and scaling using a Gaussian distribution with mean 00 and variance $l_{0,uv}^{2}(\rho)$. The proof of Theorem 3.7 is in Section C of the Supplementary Material. We next state a result regarding the asymptotic law of the quantity $\sqrt{n}\,(T_{uv}-\Psi_{0,uv})/\mathcal{S}_{0,uv}$, demonstrate how this result can be used to obtain asymptotic entrywise frequentist coverage for $\Psi_{0,uv}$ using CC-FABLE, and outline different choices of $\rho$. For general random variables $X_{n}$ and $X$, we denote $X_{n}$ converging in distribution to $X$ by $X_{n}\implies X.$
Theorem 3.8.
Suppose the Assumptions of Theorem 3.7 hold. For $1\leq u,v\leq p$, let
$\begin{split}\mathcal{S}_{0,uv}^{2}&=\left\{\begin{array}[]{ll}\sigma_{0v}^{2}\|\lambda_{0u}\|_{2}^{2}+\sigma_{0u}^{2}\|\lambda_{0v}\|_{2}^{2}+\|\lambda_{0u}\|_{2}^{2}\|\lambda_{0v}\|_{2}^{2}+(\lambda_{0u}^{\top}\lambda_{0v})^{2},&\text{for $u\neq v$},\\ 2(\|\lambda_{0u}\|_{2}^{2}+\sigma_{0u}^{2})^{2},&\text{for $u=v$.}\end{array}\right.\end{split}$
Then, as $n\to\infty$, one has $\sqrt{n}(T_{uv}-\Psi_{0,uv})/\mathcal{S}_{0,uv}\implies N(0,1)$ for all $1\leq u,v\leq p.$
The proof of Theorem 3.8 is available in Section D of the Supplementary Material. Under Theorem 3.7, the $100(1-\alpha)\%$ asymptotic pseudo-posterior credible interval of ${\Psi}_{C,uv}$ is given by
$\mathcal{C}_{uv}(\rho)=\left[T_{uv}\pm z_{1-(\alpha/2)}\dfrac{l_{0,uv}(\rho)}{\sqrt{n}}\right],$
where $z_{1-(\alpha/2)}=\Phi^{-1}\left\{1-(\alpha/2)\right\}.$ Thus, the probability of $\mathcal{C}_{uv}(\rho)$ covering $\Psi_{0,uv}$ under repeated sampling is given by
$P_{0}\left\{\Psi_{0,uv}\in\mathcal{C}_{uv}(\rho)\right\}$ $=P_{0}\left\{\dfrac{\sqrt{n}\,\left|T_{uv}-\Psi_{0,uv}\right|}{\mathcal{S}_{0,uv}}\leq z_{1-(\alpha/2)}\dfrac{l_{0,uv}(\rho)}{\mathcal{S}_{0,uv}}\right\}$
$\to q_{uv}(\rho):=2\Phi\left\{z_{1-(\alpha/2)}\dfrac{l_{0,uv}(\rho)}{\mathcal{S}_{0,uv}}\right\}-1,$
as $n\to\infty,$ using Theorem 3.8.
The form of $q_{uv}(\rho)$ suggests a general data-adaptive framework to choose $\rho$. For example, we can ensure that all the entrywise asymptotic coverages are at least $1-\alpha$ by letting $\rho=\sup_{u,v}b_{uv}$ for $1\leq u\leq v\leq p$, with $b_{uv}$ defined in Section 2.2. To see why, first let $\hat{q}_{uv}(\rho)$ be the consistent estimator of $q_{uv}(\rho)$ obtained by substituting $\sigma_{0j}^{2}$ and $\lambda_{0j}$ with $\mathcal{V}_{j}^{2}$ and $\mu_{j}$, respectively, for $j=1,\ldots,p$; these quantities are defined in Section 2. Next, observe that $\hat{q}_{uv}(\rho)$ is a non-decreasing function of $\rho>0$ for all $1\leq u,v\leq p$. It is clear that $\hat{q}_{uv}(b_{uv})=1-\alpha$ for all $1\leq u\leq v\leq p.$ Thus, $\hat{q}_{uv}\left(\sup_{u^{\prime},v^{\prime}}b_{u^{\prime}v^{\prime}}\right)\geq\hat{q}_{uv}(b_{uv})=1-\alpha.$ Alternatively, we can ensure the average of the entrywise asymptotic coverages is $1-\alpha$ by numerically solving
$\dfrac{2}{p(p+1)}\displaystyle\sum_{1\leq u\leq v\leq p}\hat{q}_{uv}(\rho)=1-\alpha$ (19)
for $\rho$. We observed in simulation experiments that simply letting
$\rho=\overline{b}=\left(\sum_{1\leq u\leq v\leq p}b_{uv}\right)/\left\{\dfrac{p(p+1)}{2}\right\}$
is a close approximation to the root of (19). In practice, we thus recommend choosing $\rho=\overline{b}$ for computational convenience, as this avoids solving the nonlinear equation. This worked well in the simulations we carried out, both from the perspective of median and average asymptotic entrywise coverages. Since Lemma E.5 in Section E of the Supplementary Material ensures that the matrix $B=(b_{uv})_{1\leq u,v\leq p}$ satisfies $\|B\|_{\infty}=O_{P_{0}}(1)$, all the aforementioned estimates of $\rho$ are bounded.
4 Simulation Results
4.1 Setup
In this Section, we compare the proposed CC-FABLE approach with competitors when judging their performance from the viewpoint of estimation error, uncertainty quantification (UQ), and computational efficiency. With $\Lambda_{0}=(\Lambda_{0,uv})_{1\leq u,v\leq p}$ the true matrix of loadings and $\Sigma_{0}=\mbox{diag}(\sigma_{01}^{2},\ldots,\sigma_{0p}^{2})$ the true matrix of idiosyncratic error variances, the data generating covariance matrix is given by $\Psi_{0}=\Lambda_{0}\Lambda_{0}^{\top}+\Sigma_{0}$. For both the estimation error and UQ experiments, we proceed as follows. We vary the number of samples $n\in\{500,1000\}$, the number of dimensions $p\in\{1000,5000\}$, and let the true number of factors be $k=10.$ For each possible $(n,p)\in\{500,1000\}\times\{1000,5000\}$, we generate $(\Lambda_{0},\Sigma_{0})$ as
$\Lambda_{0,jl}$ $\overset{ind}{\sim}0.5\,\widetilde{\delta}_{0}+0.5\,N(0,0.5^{2})$ for all $1\leq j\leq p,1\leq l\leq k$, and
$\sigma_{0j}^{2}$ $\overset{ind}{\sim}\mathcal{U}(0.5,5),$ for all $1\leq j\leq p,$
where $\widetilde{\delta}_{0}$ denotes a point mass at 00. For each combination $(n,p)$, we use the generated $\Psi_{0}=\Lambda_{0}\Lambda_{0}^{\top}+\Sigma_{0}$ and replicate the data generating process $R=100$ times. Given an estimator $\widehat{\Psi}_{0}$ of $\Psi_{0}$, we judge the estimator using the relative spectral error, defined as
$\mathcal{L}(\Psi_{0},\widehat{\Psi}_{0})=\dfrac{\|\Psi_{0}-\widehat{\Psi}_{0}\|}{\|\Psi_{0}\|},$
where $\|A\|=\sup_{\|x\|_{2}=1}\|Ax\|_{2}$. When using a Bayesian method, we consider $\widehat{\Psi}_{0}$ to be the posterior mean. For a given $(n,p)$ and a particular replicate, we obtain $\mathcal{L}(\Psi_{0},\widehat{\Psi}_{0})$ given an estimator $\widehat{\Psi}_{0}$ and then average this quantity over the replicates. For the UQ experiment, we simply consider the $95\%$ entrywise credible/pseudo-credible intervals obtained and compute the frequentist coverage over the $R=100$ replicates. To ease the computational burden, we consider the coverage of a randomly chosen $100\times 100$ submatrix of $\Psi_{0}$, corresponding to the covariance of $100$ randomly chosen variables held fixed across replicates, for each $(n,p).$ Lastly, we present a runtime analysis of the different methods considered in the simulation studies. First, we compute the theoretical computational complexity of the proposed CC-FABLE method. Next, we carry out numerical experiments to compare the computational runtimes for the different approaches.
As competitors to the proposed CC-FABLE approach, we consider the multiplicative gamma shrinkage prior formulation (Bhattacharya and Dunson,, 2011) and the approach in Ročková and George, (2016) and denote them by MGSP and ROTATE, respectively, across all the simulations. The approach in Bhattacharya and Dunson, (2011) implements posterior sampling with Markov chain Monte Carlo (MCMC), yielding posterior samples of the covariance matrix, from which we extract the posterior mean and entrywise $95\%$ credible intervals. In contrast, the approach of Ročková and George, (2016) only provides a point estimate of the covariance matrix, using expectation-maximization (EM) algorithms to approximate the posterior mean corresponding to spike and slab prior distributions on the factor loadings. As a result, we do not consider ROTATE for our UQ experiments. For CC-FABLE, we consider $1000$ Monte Carlo iterates; for estimation error experiments, we use the pseudo-posterior mean, which is analytically available. For MGSP, we consider $3000$ MCMC iterates and discard the first $1000$ as burn-in. As described earlier, CC-FABLE does not require MCMC for inference, instead providing direct Monte Carlo samples. To implement MGSP, we used the infinitefactor package (Poworoznek et al.,, 2021), while code to implement ROTATE was obtained from http://veronikarock.com/FACTOR_ANALYSIS.zip. All the simulations were carried out in the R programming language (R Core Team,, 2021).
4.2 Estimation Performance
We show the results from the simulation experiments in Table 1. For all four different combinations of $(n,p)$, the performance of CC-FABLE is either the best or very close to being the best. Both CC-FABLE and MGSP perform uniformly better than the ROTATE approach across the cases. For the most challenging case with $n=500$ and $p=5000$, CC-FABLE performs noticeably better than MGSP. We believe the well-documented issues in mixing when implementing MCMC when sampling from the posterior for MGSP to be a contributing factor. Furthermore, MGSP seems to be more affected than either CC-FABLE or ROTATE when the number of dimensions increases.
Table 1: Comparison of estimation error between CC-FABLE, MGSP, and ROTATE.
Cases CC-FABLE MGSP ROTATE
$n=500,p=1000$ 0.31 0.34 0.43
$n=1000,p=1000$ 0.22 0.21 0.28
$n=500,p=5000$ 0.32 0.40 0.43
$n=1000,p=5000$ 0.24 0.25 0.29
4.3 Frequentist Coverage
As described earlier, for each $(n,p)$, we first consider $100$ randomly chosen variables and hold them fixed across $R=100$ replicates of the data. Next, we obtain the average coverage and interval width of $95\%$ posterior/pseudo-posterior credible intervals corresponding to the relevant $100\times 100$ submatrix of $\Psi_{0}$. For all the simulation instances, we construct the entries $b_{uv}$ as described in Section 2.2, and then proceed to set the coverage-correction factor $\rho$ to be the average of $\{b_{uv}:1\leq u\leq v\leq p\}$ for coverage-correction. The results, provided in Table 2, suggest that CC-FABLE provides superior entrywise coverage when compared to MGSP, with comparable interval width. In all the cases, the average coverage of entrywise intervals obtained from MGSP falls short of the nominal value $0.95$, while CC-FABLE meets the nominal coverage. Furthermore, the average entrywise coverage obtained from MGSP decreases when the number of dimensions increases for a fixed sample size, while the results of CC-FABLE are not affected.
Table 2: Comparison of average frequentist coverages and average interval widths across entrywise pseudo-credible intervals obtained from CC-FABLE and MGSP.
Cases Coverage Coverage Width Width
$(n,p)$ (CC-FABLE) (MGSP) (CC-FABLE) (MGSP)
$n=500,p=1000$ 0.95 0.84 0.48 0.48
$n=1000,p=1000$ 0.95 0.72 0.35 0.31
$n=500,p=5000$ 0.95 0.81 0.50 0.46
$n=1000,p=5000$ 0.95 0.69 0.35 0.31
4.4 Computational Efficiency
For the purposes of computing order complexity, we assume $k$ is fixed and $n,p$ vary. We first consider the order complexity when generating $N_{0}$ pseudo-posterior samples of the covariance matrix. The first stage SVD of the $n\times p$ matrix $\mathbf{Y}$ has $\mathcal{O}(np\min(n,p))$ complexity (Vasudevan and Ramakrishna,, 2017), which needs to be only computed once. Computing the $p$ hyperparameters $\mu_{j}$ and $\delta_{j}^{2}$ for $j=1,\ldots,p$ has $\mathcal{O}(np)$ complexity, while generating each Monte Carlo sample of $({\Lambda}_{C},{\Sigma}_{C})$ for $j=1,\ldots,p$ has $\mathcal{O}(p)$ complexity since $({\lambda}_{C,j},{\sigma}_{C,j}^{2})$ are independently generated for $j=1,\ldots,p$. Lastly, obtaining $\Psi_{C}=\Lambda_{C}\Lambda_{C}^{\top}+\Sigma_{C}$ has $\mathcal{O}(p^{2})$ complexity for each sample, leading to $\mathcal{O}(p^{2}N_{0})$ complexity to obtain $N_{0}$ samples. Thus, obtaining $N_{0}$ pseudo-posterior samples of the covariance matrix after coverage-correction has $\mathcal{O}\{p^{2}N_{0}+pN_{0}+np\min(n,p)\}=\mathcal{O}\{p^{2}N_{0}+np\min(n,p)\}$ complexity. Next, we consider the order complexity to obtain the pseudo-posterior mean of the covariance. As before, we would require computing the SVD with $\mathcal{O}\{np\min(n,p)\}$ complexity, while computing the hyperparameters $(\mu_{j},\delta_{j}^{2})$ for $j=1,\ldots,p$ has $\mathcal{O}(np)$ complexity. Obtaining the matrix $G=G_{0}G_{0}^{\top}$ from $G_{0}$ has $\mathcal{O}(p^{2})$ complexity. Thus, computing the pseudo-posterior mean of $\Psi$ has asymptotic complexity $\mathcal{O}\{p^{2}+np\min(n,p)\}$.
The runtime for both sampling and pseudo-posterior mean computation may be substantially improved by considering computationally efficient implementations of the SVD step such as the randomized SVD (Halko et al.,, 2011) when either $n$ or $p$ is very large. Additionally, since the CC-FABLE algorithm employs embarrassingly parallel sampling, the computational runtime would be further improved by parallelization of the computations across cores.
In Figure 1, we compare the three methods CC-FABLE, MGSP, and ROTATE in terms of obtaining (pseudo) posterior samples and point estimates of the covariance matrix. We consider the same data-generating scenario as in Section 4.1. We fix $n=500$ and vary $p$ from $500$ to $5000$ in increments of $500$, obtaining runtime measurements across $R=20$ independent replicates for each choice of $p$. As before, we consider $3000$ MCMC iterations for MGSP and discard the first $1000$ iterates as burn-in. The experiments were carried out on an M1 MacBook Pro with 32 GB of RAM (random-access memory). We compare CC-FABLE with MGSP when obtaining samples of the factor loadings and the error variances, and with ROTATE when obtaining a point estimate of the covariance matrix. In both cases, CC-FABLE is faster compared to its competitors. When obtaining samples, CC-FABLE is about $150-200$ times faster than MGSP, with this ratio remaining fairly constant as $p$ increases. When obtaining a point estimator of the covariance, CC-FABLE is at least twice as fast as ROTATE.
Refer to caption Refer to caption
Figure 1: Plots comparing the runtimes in seconds (in $\log_{10}$ scale) of CC-FABLE, MGSP, and ROTATE approaches. The first plot compares the runtimes of CC-FABLE and MGSP when obtaining (pseudo) posterior samples of the covariance matrix. The second plot compares runtimes of CC-FABLE and ROTATE when obtaining a point estimate of the covariance matrix.
5 Application
In this Section, we apply our method to identify associations between expressed genes in a high-throughput sequencing immune cell dataset. Identifying gene expression associations in varying immune cell populations and among multiple lineages is critical to understanding the inflammatory response to infectious and chronic diseases (Ota et al.,, 2021; Saini et al.,, 2022). This enables subsequent detection of corresponding cellular and molecular pathways underlying pathological conditions. For this application, we consider the GSE109125 dataset, which contains RNAseq data from 127 highly purified immune cell lineages including adaptive and innate lymphocytes, myeloid cells, mast cells, and neutrophils. The lineages along with the associated differentiation cascades were profiled using the ImmGen ULI pipeline (Yoshida et al.,, 2019).
We first log-transformed raw data counts $R_{ij}$ for the $i$th cell and $j$th gene as $y^{*}_{ij}=\log_{2}(R_{ij}+1)$. Next, we proceed to center the data for analysis before fitting the latent factor model. Finally, we use the genefilter package in R (Gentleman et al.,, 2023) to filter relevant genes. In this case, we filtered genes according to their variance and considered the top $10\%$ of genes with the highest variances. There were no cells with missing entries. After carrying out appropriate pre-processing of the dataset, we obtain expression data $\mathbf{Y}\in\mathbb{R}^{n\times p}$ for $p=5300$ genes measured on $n=205$ cells.
We compare results from applying the three methods CC-FABLE, MGSP, and ROTATE on the data. The fitted log-likelihoods using the whole data $\mathbf{Y}$ are $-490006.4$, $-503574$, and $-728959.9$ for CC-FABLE, MGSP, and ROTATE, respectively. This suggests that CC-FABLE fits the data the best among the three methods. The estimated rank when using the $\mbox{JIC}(k)$ criterion as highlighted in Section 2.3 is $\widehat{k}=30$. As a posterior predictive check, we investigated the average coverage of entrywise predictive intervals across all the entries of $\mathbf{Y}$. When using CC-FABLE, the average predictive coverage is $0.97$. To further investigate model fit, we also looked into the proportion of variance explained for each gene using CC-FABLE. The average proportion of variance explained across variables when using CC-FABLE is $0.83$, indicating that the factor model is a good fit.
We now focus on estimating a $100\times 100$ submatrix of the $5300\times 5300$ covariance matrix, corresponding to the $100$ genes with the highest variability. Suppose the indices of these genes are $\mathcal{I}_{0}\subset\{1,2,\ldots,p\}$. We assess estimation performance using out-of-sample fit on a randomly chosen $50\times 100$ submatrix $\mathbf{Y}_{T}$ of $\mathbf{Y}$, corresponding to expression data on the $100$ genes in $\mathcal{I}_{0}$ for $50$ cells. To estimate the covariance of genes in $\mathcal{I}_{0}$, it is typical to only use expression data for those genes. Alternatively, we infer the covariance of these genes jointly with additional genes with indices in the set $\mathcal{A}(p_{S})$ of length $p_{S}>0$ such that $\mathcal{A}(p_{S})\cap\mathcal{I}_{0}=\phi$. We vary $p_{S}\in\{100j:j=1,\ldots,10\}\cup\{2000,4000\}$ and extract the relevant $100\times 100$ covariance $\widehat{\Psi}(p_{S})$ of the genes with indices in $\mathcal{I}_{0}$. We then use $\widehat{\Psi}(p_{S})$ to evaluate the out-of-sample log-likelihood (OOSLL) on the held-out test set $\mathbf{Y}_{T}$. We carry out this analysis for $R=10$ independent replicates of the train-test split and compare the methods CC-FABLE, MGSP, and ROTATE on the basis of average OOSLL. The results are illustrated in Figure 2.
In Figure 2, it is clear that both CC-FABLE and MGSP exhibit superior performance to ROTATE. Both MGSP and ROTATE exhibit an essentially monotonically decreasing trend as extra genes are added, so including additional genes does not improve inferences on the genes of interest in using these methods. In sharp contrast, CC-FABLE has rapidly improved performance initially as extra genes are added, exhibiting a clear blessing of dimensionality. After a few hundred genes, the gain levels off, and eventually there is a modest decline, perhaps due to the need to add additional factors not related to the genes of interest when large numbers of additional genes are added.
The proposed method thus shows promising results in this high-dimensional gene expression data application and compares favourably to the two state-of-the-art gold standard approaches to Bayesian factor analysis. The results obtained from CC-FABLE come at a fraction of the computational budget compared to MGSP due to its embarrassingly parallel sampling scheme requiring no MCMC. When fitted on the full dataset on an M1 MacBook Pro with $32$ GB of RAM, CC-FABLE only took $1.1$ seconds to compute the pseudo-posterior mean while running MGSP with $3000$ MCMC and $1000$ burn-in iterates took $\sim 27$ minutes. This translates to a speed-up of close to $1600$ times when using CC-FABLE.
Figure 2: Figure showing log-likelihood of the test set (y-axis) averaged over $R=10$ replicates of the train-test split, when using additional genes (x-axis) to estimate the covariance of the relevant set of $100$ variables, using CC-FABLE, MGSP, and ROTATE. The training set has $155$ cells and the test set has $50$ cells for each of the $R=10$ train-test split replicates.
6 Discussion
In this paper, we developed a computationally scalable approach to fit high-dimensional Bayesian latent factor models with Gaussian data. Leveraging on a blessing of dimensions, the proposed FABLE approach completely bypasses the need to carry out MCMC to provide pseudo-posterior samples of the covariance matrix for inference. By avoiding MCMC altogether, our approach averts the pitfalls of MCMC in higher dimensions, such as slow mixing and unsatisfactory convergence. Due to its embarrassingly parallel nature, the FABLE approach leads to immense computational benefits over current approaches dependent on MCMC. We developed a coverage-correction modification to FABLE called CC-FABLE with provable guarantees on asymptotic frequentist coverage of its pseudo-credible intervals. The CC-FABLE pseudo-posterior also enjoys desirable theoretical properties such as consistency and compares favorably to state-of-the-art approaches to Bayesian factor analysis from the viewpoint of estimation error, uncertainty quantification, and computational efficiency in our numerical experiments and gene expression data application.
This work naturally leads to several important directions for future research. The first direction is to model non-Gaussian data, such as binary or count data. Such data are routinely collected in ecological studies. For example, $y_{ij}$ for $i=1,\ldots,n$ and $j=1,\ldots,p$ could indicate the presence/absence or abundance of species $j$ in sample $i$. For presence/absence data, let $\mbox{P}(y_{ij}=1\mid\alpha_{0},\eta_{i},\lambda_{j})=\mbox{F}\left(\alpha_{0}+\eta_{i}^{\top}\lambda_{j}\right),$ with $\eta_{i}\overset{iid}{\sim}N_{k}(0,\mathbb{I}_{k})$. Here, $F$ is a link function, such as the probit or the logistic link. The observations $y_{ij}$ are mutually independent conditional on the latent factors $\eta_{i}$, but integrating them out induces dependence between $y_{i1},\ldots,y_{ip}.$ Bayesian inference in such models typically proceeds via a data augmentation Gibbs sampler by introducing auxiliary continuous variables $z_{ij}$ and alternating between sampling $\eta_{i},\lambda_{j},$ and $z_{ij}$ for $i=1,\ldots,n$ and $j=1,\ldots,p$ given the data $\mathbf{Y}=(y_{ij})$ using MCMC (Albert and Chib,, 1993; Polson et al.,, 2013). However, in our motivating ecological settings involving insects and fungi, $p$ is on the order of 10,000100,$0-100,00$ 100 , 000, rendering current MCMC-based Bayesian factor analytic approaches computationally infeasible. To extend FABLE to this setting, we can pre-estimate the latent factors via an SVD or an appropriate generalized linear model (GLM) variant, and then fit the resulting Bayesian GLMs in parallel for each species, potentially relying on Laplace approximations.
An additional important direction is to modify the FABLE inferential framework to allow more intricate hierarchical modeling. For example in Bayesian factor analysis, it is often of interest to (1) include covariate effects, (2) use more elaborate priors on the loadings to favor sparsity and adaptive selection of the number of factors, (3) consider more flexible latent factor distributions, and/or (4) allow nonlinear measurement structure. In exploratory analyses, we have observed good performance for ad hoc FABLE modifications to include covariates and for simulated data having non-Gaussian latent factor distributions. Problem (2) can potentially be addressed in the second stage via MCMC algorithms implemented in parallel for the different sparse regressions after inferring the latent factors using a sparse SVD in a first stage. Problem (4) may necessitate non-linear dimensionality reduction in the first stage. Obtaining concrete implementations with theory support for problems (1)-(4) is of future interest.
The current work also hints at further investigation into the theoretical properties of Bayesian factor models. Although there are existing results focusing on posterior consistency with carefully chosen shrinkage priors on the factor loadings, these are primarily in the case where the idiosyncratic model errors are Gaussian. It would thus be interesting to explore how the blessing of dimensions influences both posterior consistency and uncertainty quantification in different settings, such as when the data are non-Gaussian. Furthermore, there remains a clear need to extend existing theoretical results on Bayesian factor models to results on uncertainty quantification of the covariance matrix.
Acknowledgments
This work was partially supported by Merck, the United States Office of Naval Research (ONR) under grant N00014-21-1-2510, the European Research Council (ERC) under grant 856506, and the National Institutes of Health (NIH) under grant R01ES035625. Code used to implement CC-FABLE used in the simulation experiments and the gene expression data application is available to download at https://github.com/shounakch/FABLE.
Data Availability
The gene expression data set used in Section 5 of this paper is publicly available at https://www.ncbi.nlm.nih.gov/geo/query/acc.cgi?acc=GSE109125.
Supplementary Material
Details on the content of this supplement are as follows.
  1. 1.
    Section A contains the proof of Proposition 2.1 in the main paper.
  2. 2.
    Section B contains the proof of Theorem 3.6 in the main paper.
  3. 3.
    Section C contains the proof of Theorem 3.7 in the main paper.
  4. 4.
    Section D contains the proof of Theorem 3.8 in the main paper.
  5. 5.
    Section E contains the proof of Proposition 3.5 and relevant lemmas used to prove Theorem 3.6 in the main paper.
  6. 6.
    Section F contains the proof of relevant lemmas used to prove Theorems 3.7 and 3.8 in the main paper.
Appendix A Proof of Proposition 2.1
Proof.
  1. 1.
    We first have $\widehat{\mathbf{M}}^{\top}\widehat{\mathbf{M}}=\widehat{C}^{-1}\mathbf{A}^{\top}\mathbf{A}(\widehat{C}^{\top})^{-1}=\widehat{C}^{-1}(n\widehat{C}\widehat{C}^{\top})(\widehat{C}^{\top})^{-1}=n\mathbb{I}_{k}$. Next, we have $\widehat{\mathbf{M}}\widehat{\mathbf{M}}^{\top}=\mathbf{A}(\widehat{C}\widehat{C}^{\top})^{-1}\mathbf{A}^{\top}=(UD/\sqrt{p})(npD^{-2})(DU^{\top}/\sqrt{p})=nUU^{\top}$.
  2. 2.
    For any $\widehat{\mathbf{M}}$ satisfying the condition, $\mathbf{K}=\{\widehat{\mathbf{M}}^{\top}\widehat{\mathbf{M}}+(\mathbb{I}_{k}/\tau^{2})\}^{-1}=(n+\tau^{-2})^{-1}\mathbb{I}_{k}.$ It is clear that the pseudo-posterior of $\widetilde{\sigma}_{j}^{2}$ only depends on $\widehat{\mathbf{M}}$ through $\widehat{\mathbf{M}}^{\top}\widehat{\mathbf{M}}$ and $\widehat{\mathbf{M}}\widehat{\mathbf{M}}^{\top}$ for all $1\leq j\leq p$. For $1\leq u\leq p,$ a pseudo-posterior draw of $\tilde{\lambda}_{u}$ may be represented as
    $\widetilde{\lambda}_{u}=\mu_{u}+\widetilde{\sigma}_{u}\mathbf{K}^{1/2}Z_{u},$
    where $Z_{u}\overset{ind}{\sim}N_{k}(0_{k},\mathbb{I}_{k})$ for $1\leq u\leq p$ and $Z_{u}$ is independent of $\widetilde{\sigma}_{u}^{2}$. This implies that for $1\leq u,v\leq p$,
    $\widetilde{\lambda}_{u}^{\top}\widetilde{\lambda}_{v}=\mu_{u}^{\top}\mu_{v}+\widetilde{\sigma}_{v}\mu_{u}^{\top}\mathbf{K}^{1/2}Z_{v}+\widetilde{\sigma}_{u}\mu_{v}^{\top}\mathbf{K}^{1/2}Z_{u}+\widetilde{\sigma}_{u}\widetilde{\sigma}_{v}Z_{u}^{\top}\mathbf{K}Z_{v}.$
    We now consider each expression one-by-one. The first term is $\mu_{u}^{\top}\mu_{v}=y^{(u)\top}\widehat{\mathbf{M}}\mathbf{K}^{2}\widehat{\mathbf{M}}^{\top}y^{(v)}=(n+\tau^{-2})^{-2}y^{(u)\top}\widehat{\mathbf{M}}\widehat{\mathbf{M}}^{\top}y^{(v)}.$ When $u\neq v$, the second and third terms have Gaussian distributions with means 00 and variances $(n+\tau^{-2})^{-1}\widetilde{\sigma}_{u}^{2}\|\mu_{v}\|_{2}^{2}$ and $(n+\tau^{-2})^{-1}\widetilde{\sigma}_{v}^{2}\|\mu_{u}\|_{2}^{2}$, respectively. When $u=v$, both terms are equal and the sum has a Gaussian distribution with mean 00 and variance $4\widetilde{\sigma}_{u}^{2}\|\mu_{u}\|_{2}^{2}$. The final term is equal to $(n+\tau^{-2})^{-1}\widetilde{\sigma}_{u}\widetilde{\sigma}_{v}Z_{u}^{\top}Z_{v}$. Thus, all the terms depend on $\widehat{\mathbf{M}}$ only as hypothesized, implying the result for $\widetilde{L}$. The result for $\widetilde{\Psi}$ is obtained by simply observing that $\widetilde{\Psi}=\widetilde{L}+\widetilde{\Sigma}.$
This proves the desired result. ∎
Appendix B Proof of Theorem 3.6
B.1 Proof of part (a)
Proof.
We first start with the pseudo-posterior contraction of $L_{C}$ to $L_{0}.$ Let $G_{0}=\sqrt{n}\mathbf{Y}^{\top}U/(n+\tau^{-2})$ and $G=G_{0}G_{0}^{\top}=n\mathbf{Y}^{\top}UU^{\top}\mathbf{Y}/(n+\tau^{-2})^{2}.$ We have
$\begin{split}L_{C}-L_{0}&={\Lambda}_{C}{\Lambda}_{C}^{\top}-\Lambda_{0}\Lambda_{0}^{\top}\\ &=\left[\frac{\mathbf{Y}^{\top}\widehat{M}\widetilde{E}^{\top}+\widetilde{E}\widehat{M}^{\top}\mathbf{Y}}{n+\tau^{-2}}+\widetilde{E}\widetilde{E}^{\top}\right]+\left(G-\Lambda_{0}\Lambda_{0}^{\top}\right),\end{split}$
by expressing $\Lambda_{C}=G_{0}+\widetilde{E}$, where $\widetilde{E}=[\tilde{e}_{1},\ldots,\tilde{e}_{p}]^{\top}$ with $\tilde{e}_{j}\overset{ind}{\sim}N_{k}\left(0,\frac{\rho^{2}\sigma_{C,j}^{2}}{n+\tau^{-2}}\mathbb{I}_{k}\right)$ for $j=1,\ldots,p$. We decompose
$\begin{split}&G-\Lambda_{0}\Lambda_{0}^{\top}=\frac{n}{(n+\tau^{-2})^{2}}\left(\mathbf{Y}^{\top}\mathbf{Y}-\mathbf{Y}^{\top}U_{\perp}U_{\perp}^{\top}\mathbf{Y}\right)-\Lambda_{0}\Lambda_{0}^{\top}\\ &=\frac{-n}{(n+\tau^{-2})^{2}}\mathbf{Y}^{\top}U_{\perp}U_{\perp}^{\top}\mathbf{Y}+\frac{n}{(n+\tau^{-2})^{2}}\left(M_{0}\Lambda_{0}^{\top}+E\right)^{\top}\left(M_{0}\Lambda_{0}^{\top}+E\right)-\Lambda_{0}\Lambda_{0}^{\top}\\ &=\frac{-n}{(n+\tau^{-2})^{2}}\mathbf{Y}^{\top}U_{\perp}U_{\perp}^{\top}\mathbf{Y}+\frac{n}{(n+\tau^{-2})^{2}}\Lambda_{0}(M_{0}^{\top}M_{0}-nI)\Lambda_{0}^{\top}+\left(\frac{n^{2}}{(n+\tau^{-2})^{2}}-1\right)\Lambda_{0}\Lambda_{0}^{\top}+\\ &\frac{n}{(n+\tau^{-2})^{2}}(\Lambda_{0}M_{0}^{\top}E+E^{\top}M_{0}\Lambda_{0}^{\top}+E^{\top}E).\end{split}$
To develop an upper bound of $\|L_{C}-L_{0}\|$, we aim to develop an upper bound for the spectral norm of each term in the inequality above. We enumerate them as follows.
  • (i)
    First, we develop a probabilistic upper bound for ${\sigma}_{C,j}^{2}.$ Recall that $\|y^{(j)}\|_{2}^{2}=\sum_{i=1}^{n}y_{ij}^{2}\sim(\sigma_{0j}^{2}+\|\lambda_{0j}\|_{2}^{2})\,\chi_{n}^{2}.$ We apply the tail inequality bound of the $\chi^{2}$ distribution following Lemma 1 in Laurent and Massart, (2000) to obtain $\|y^{(j)}\|_{2}\lesssim n$ with probability at least $1-o(1)$. Thus, with probability at least $1-o(1)$ we have
    $\frac{\gamma_{0}\delta_{0}^{2}+\|y^{(j)}\|_{2}^{2}-(n+\tau^{-2})\|\mu_{j}\|_{2}^{2}}{2}\lesssim n.$
    As
    ${\sigma}_{C,j}^{2}\sim\mbox{IG}\left(\frac{\gamma_{0}+n}{2},\frac{\gamma_{0}\delta_{0}^{2}+\|y^{(j)}\|_{2}^{2}-(n+\tau^{-2})\|\mu_{j}\|_{2}^{2}}{2}\right),$
    the Gamma distribution tail bound in Theorem 5 of Zhang and Zhou, (2020) implies there exists a constant $C^{\prime}>0$ such that
    $\begin{split}\Pi_{C}\left(\max_{1\leq j\leq p}{\sigma}_{C,j}^{2}\leq C^{\prime}\right)=1-o_{P_{0}}(1).\end{split}$ (20)
  • (ii)
    Since $\|\Lambda_{0}\|\asymp\sqrt{p_{n}}$, we have
    $\left\|\left\{\frac{n^{2}}{(n+\tau^{-2})^{2}}-1\right\}\Lambda_{0}\Lambda_{0}^{\top}\right\|=\left|\frac{n^{2}-(n+\tau^{-2})^{2}}{(n+\tau^{-2})^{2}}\right|\|\Lambda_{0}\|^{2}\lesssim\dfrac{p_{n}}{n}.$ (21)
  • (iii)
    Recall $M_{0}$ is an $n$-by-$k$ matrix with i.i.d. Gaussian entries. By Lemma E.1, we have with probability at least $1-o(1)$ the following:
    $\left\|\frac{n}{(n+\tau^{-2})^{2}}\Lambda_{0}(M_{0}^{\top}M_{0}-nI)\Lambda_{0}^{\top}\right\|$ $\lesssim\frac{1}{n}\|\Lambda_{0}\|^{2}\|M_{0}^{\top}M_{0}-nI\|$
    $\lesssim\frac{1}{n}\|\Lambda_{0}\|^{2}\|M_{0}^{\top}M_{0}-nI\|$
    $\lesssim\frac{p_{n}(k+\sqrt{nk})}{n}\lesssim\dfrac{p_{n}}{\sqrt{n}}.$
  • (iv)
    Next, we look at
    $\begin{split}&\left\|\frac{n}{(n+\tau^{-2})^{2}}(\Lambda_{0}^{\top}M_{0}E+E^{\top}M_{0}\Lambda_{0}^{\top}+E^{\top}E)\right\|\lesssim\frac{1}{n}\left(2\left\|\Lambda_{0}^{\top}M_{0}E\right\|+\left\|E^{\top}E\right\|\right).\end{split}$
    By Lemma E.2, $\|E^{\top}E\|=\|E\|^{2}\lesssim n+p_{n}+\sqrt{np_{n}}$ with probability at least $1-o(1)$ and $\|\Lambda_{0}M_{0}^{\top}E\|\leq\|\Lambda_{0}\|\,\|M_{0}\|\,\|E\|\lesssim\sqrt{np_{n}}(\sqrt{n}+\sqrt{p_{n}})$ with probability at least $1-o(1)$. Thus, with probability at least $1-o(1)$,
    $\left\|\frac{n}{(n+\tau^{-2})^{2}}(\Lambda_{0}^{\top}M_{0}E+E^{\top}M_{0}\Lambda_{0}^{\top}+E^{\top}E)\right\|\lesssim\frac{n+p_{n}+p_{n}\sqrt{n}+n\sqrt{p_{n}}+\sqrt{np_{n}}}{n}\lesssim\dfrac{p_{n}\sqrt{n}+n\sqrt{p_{n}}}{n}.$
  • (v)
    Since $U$ corresponds to the top $k$ left singular vectors of $\mathbf{Y}$, we have $\|\mathbf{Y}^{\top}U_{\perp}U_{\perp}^{\top}\mathbf{Y}\|=\|\mathbf{Y}^{\top}U_{\perp}\|^{2}=s_{k+1}^{2}(\mathbf{Y})$, where $s_{k+1}^{2}(\mathbf{Y})$ is the $(k+1)$-th singular value of $\mathbf{Y}$. By Eckart–Young Theorem (Lemma E.3), we have
    $\|\mathbf{Y}^{\top}U_{\perp}U_{\perp}^{\top}\mathbf{Y}\|=s_{k+1}^{2}(\mathbf{Y})=\min_{B:\text{rank}(B)\leq k}\|\mathbf{Y}-B\|^{2}\leq\|\mathbf{Y}-M_{0}\Lambda_{0}^{\top}\|^{2}=\|E\|^{2}\lesssim n+p_{n}+\sqrt{np_{n}}.$
    Then,
    $\left\|\frac{-n}{(n+\tau^{-2})^{2}}\mathbf{Y}^{\top}U_{\perp}U_{\perp}^{\top}\mathbf{Y}\right\|\lesssim\frac{n+p_{n}+\sqrt{np_{n}}}{n}.$
  • (vi)
    Finally, we consider
    $\begin{split}\left\|\frac{\mathbf{Y}^{\top}\widehat{M}\widetilde{E}^{\top}+\widetilde{E}\widehat{M}^{\top}\mathbf{Y}}{n+\tau^{-2}}+\widetilde{E}\widetilde{E}^{\top}\right\|\lesssim\dfrac{2}{n}\left\|\mathbf{Y}^{\top}\widehat{M}\widetilde{E}^{\top}\right\|+\|\widetilde{E}\widetilde{E}^{\top}\|.\end{split}$
    By Lemma E.2, the following hold with probability at least $1-o_{P_{0}}(1)$:
    $\|\widetilde{E}\widetilde{E}^{\top}\|=\|\widetilde{E}\|^{2}\lesssim\frac{p_{n}}{n}\rho^{2}\left(\max_{1\leq j\leq p}{\sigma}_{C,j}^{2}\right)\lesssim\frac{p_{n}}{n}.$
    $\frac{1}{n}\left\|\mathbf{Y}^{\top}\widehat{M}\widetilde{E}^{\top}\right\|$ $\lesssim\frac{1}{n}\|\mathbf{Y}^{\top}\sqrt{n}U\|\left\|\widetilde{E}\right\|\lesssim\frac{\rho}{\sqrt{n}}\|\mathbf{Y}\|\sqrt{\frac{p_{n}}{n}\left(\max_{1\leq j\leq p}{\sigma}^{2}_{C,j}\right)}$
    $\lesssim\left(\frac{\sqrt{np_{n}}+\sqrt{n}+\sqrt{p_{n}}}{\sqrt{n}}\right)\sqrt{\frac{p_{n}}{n}}\lesssim\frac{p_{n}}{\sqrt{n}}.$
    Thus, with probability at least $1-o_{P_{0}}(1)$, we have
    $\left\|\frac{\mathbf{Y}^{\top}\widehat{M}\widetilde{E}^{\top}+\widetilde{E}\widehat{M}^{\top}\mathbf{Y}}{(n+\tau^{-2})}+\widetilde{E}\widetilde{E}^{\top}\right\|\lesssim\dfrac{p_{n}}{\sqrt{n}}.$
By combining the previous steps (i)-(vi) and $\|\Lambda_{0}\|\asymp\sqrt{p_{n}}$, we have proved the desired result. ∎
B.2 Proof of part (b)
Proof.
We now show the contraction result for the pseudo-posterior of $\Sigma_{C}$. We start out by observing that $||{\Sigma}_{C}-\Sigma_{0}||=\max_{1\leq j\leq p}|D_{j}|$, where $D_{j}={\sigma}_{C,j}^{2}-\sigma_{0j}^{2}$ for $j=1,\ldots,p.$ Let $\kappa_{j}={\sigma}_{C,j}^{-2}$, so that $\kappa_{j}\overset{ind}{\sim}G(\gamma_{n}/2,\gamma_{n}\delta_{j}^{2}/2)$ under $\Pi_{C}$. Let
$U_{j}=\dfrac{\gamma_{n}\delta_{j}^{2}}{2}\kappa_{j}-\dfrac{\gamma_{n}}{2}.$
We can now express $D_{j}$ as
$D_{j}=\left(1+\dfrac{2U_{j}}{\gamma_{n}}\right)^{-1}\,\left\{(\delta_{j}^{2}-\sigma_{0j}^{2})-\dfrac{2U_{j}}{\gamma_{n}}\sigma_{0j}^{2}\right\}.$
Using Lemma E.7, we have $\max_{1\leq j\leq p}|U_{j}|/\gamma_{n}\lesssim\{(\log p_{n})/n\}^{1/3}$ with probability at least $1-o_{P_{0}}(1).$ Thus, $\min_{1\leq j\leq p}|1+(2U_{j}/\gamma_{n})|\gtrsim 1/2$ with probability at least $1-o_{P_{0}}(1).$ Therefore with probability at least $1-o_{P_{0}}(1),$
$\max_{1\leq j\leq p}|D_{j}|\lesssim 2\max_{1\leq j\leq p}|\delta_{j}^{2}-\sigma_{0j}^{2}|+4\sigma_{0j}^{2}\max_{1\leq j\leq p}\dfrac{|U_{j}|}{\gamma_{n}}.$
Let $Q_{j}=\delta_{j}^{2}-\sigma_{0j}^{2}$. Using Lemma E.6, we can represent
$Q_{j}=\dfrac{\sigma_{0j}^{2}}{n}\left\{\dfrac{Z_{j}}{\sigma_{0j}^{2}}-(n-k)\right\}-\dfrac{k\sigma_{0j}^{2}}{n}+F_{j},$
where $Z_{j}/\sigma_{0j}^{2}\sim\chi_{n-k}^{2}\equiv G\{(n-k)/2,1/2\}$ and $\max_{1\leq j\leq p}|F_{j}|\lesssim\dfrac{1}{n}+\dfrac{1}{p_{n}}$ with probability at least $1-o(1).$ Using Lemma E.7, we obtain
$\max_{1\leq j\leq p}\left|\dfrac{Z_{j}}{\sigma_{0j}^{2}}-(n-k)\right|\lesssim n\left(\dfrac{\log p_{n}}{n}\right)^{1/3}$
with probability at least $1-o(1)$. Thus, with probability at least $1-o(1)$, we get
$\max_{1\leq j\leq p}|Q_{j}|\lesssim\left(\dfrac{\log p_{n}}{n}\right)^{1/3}+\dfrac{1}{p_{n}}.$
From Lemma E.7, we have $\max_{1\leq j\leq p}|U_{j}|/\gamma_{n}\lesssim\left({\log p_{n}}/{n}\right)^{1/3}$ with probability at least $1-o_{P_{0}}(1)$. Thus, we obtain
$\max_{1\leq j\leq p}|D_{j}|\lesssim\left(\dfrac{\log p_{n}}{n}\right)^{1/3}+\dfrac{1}{p_{n}}$
with probability at least $1-o_{P_{0}}(1),$ for all $j=1,\ldots,p$ or equivalently,
$\|{\Sigma}_{C}-\Sigma_{0}\|\lesssim\left(\dfrac{\log p_{n}}{n}\right)^{1/3}+\dfrac{1}{p_{n}}$
with probability at least $1-o_{P_{0}}(1)$. This proves the result. ∎
B.3 Proof of part (c)
Proof.
Under the conditions of Theorem 1, we have $\|\Psi_{0}\|\asymp\|L_{0}\|\asymp p_{n}.$ By the triangle inequality, $\|\Psi_{C}-\Psi_{0}\|\leq\|L_{C}-L_{0}\|+\|\Sigma_{C}-\Sigma_{0}\|$. Thus, we have under $\Pi_{C}$ that
$\dfrac{\|{\Psi}_{C}-\Psi_{0}\|}{\|\Psi_{0}\|}\lesssim\dfrac{1}{\sqrt{p_{n}}}+\dfrac{1}{\sqrt{n}}+o\left(\dfrac{1}{p_{n}}\right)\lesssim\dfrac{1}{\sqrt{p_{n}}}+\dfrac{1}{\sqrt{n}}$
with probability at least $1-o_{P_{0}}(1)$, proving the result. ∎
Appendix C Proof of Theorem 3.7
Proof.
First let $u\neq v$, so that $\Psi_{C,uv}=L_{C,uv}$, where $L_{C,uv}$ is the $uv$th element of $L_{C}=\Lambda_{C}\Lambda_{C}^{\top}$. According to the pseudo-posterior generation mechanism after coverage correction,
$L_{C,uv}=\mu_{u}^{\top}\mu_{v}+\dfrac{\rho\left({\sigma}_{C,v}\mu_{u}^{\top}\tilde{e}_{C,v}+{\sigma}_{C,u}\mu_{v}^{\top}\tilde{e}_{C,u}\right)}{\sqrt{n+\tau^{-2}}}+\dfrac{\rho^{2}{\sigma}_{C,u}{\sigma}_{C,v}\tilde{e}_{C,u}^{\top}\tilde{e}_{C,v}}{n+\tau^{-2}},$
for $\widetilde{e}_{C,u},\widetilde{e}_{C,v}\overset{ind}{\sim}N_{k}(0,\mathbb{I}_{k})$. We can express
$\rho\left(\sigma_{C,v}\mu_{u}^{\top}\tilde{e}_{C,v}+\sigma_{C,u}\mu_{v}^{\top}\tilde{e}_{C,u}\right)$ $=l_{0,uv}(\rho)R_{uv}+d_{uv}(\rho)R_{uv},$
for $R_{uv}\sim N(0,1)$ such that $R_{uv}$ is independent of $\mathbf{Y},\sigma_{C,u}^{2},$ and $\sigma_{C,v}^{2}$, with $l_{0,uv}^{2}(\rho)=\rho^{2}\left(\sigma_{0v}^{2}\|\lambda_{0u}\|_{2}^{2}+\sigma_{0u}^{2}\|\lambda_{0v}\|_{2}^{2}\right)$ and $d_{uv}(\rho)=\rho(\sigma_{C,v}^{2}||\mu_{u}||_{2}^{2}+\sigma_{C,u}^{2}||\mu_{v}||_{2}^{2})^{1/2}-l_{0,uv}(\rho)$. We can now express
$\sqrt{n}(L_{C,uv}-\mu_{u}^{\top}\mu_{v})=\dfrac{\sqrt{n}}{\sqrt{n+\tau^{-2}}}l_{0,uv}(\rho)R_{uv}+d_{uv}^{*}(\rho),$
where
$d_{uv}^{*}(\rho)=\sqrt{\dfrac{n}{n+\tau^{-2}}}d_{uv}(\rho)R_{uv}+\sqrt{n}\left(\dfrac{\rho^{2}\sigma_{C,v}\sigma_{C,u}\tilde{e}_{C,u}^{\top}\tilde{e}_{C,v}}{n+\tau^{-2}}\right).$
Due to pseudo-posterior concentration of $\sigma_{C,j}^{2}$ to $\sigma_{0j}^{2}$ and convergence in probability of $||\mu_{j}||_{2}^{2}$ to $||\lambda_{0j}||_{2}^{2}$, we have pseudo-posterior concentration of $d_{uv}(\rho)$ around 00 as $n\to\infty$ for any finite $\rho>0$. Furthermore, with probability at least $1-o_{P_{0}}(1)$,
$\sqrt{n}\left(\dfrac{{\sigma}_{C,u}{\sigma}_{C,v}\tilde{e}_{C,u}^{\top}\tilde{e}_{C,v}}{n+\tau^{-2}}\right)\lesssim\dfrac{1}{\sqrt{n}}.$
Thus, the pseudo-posterior of $d_{uv}^{*}(\rho)$ concentrates around 00 asymptotically. As $n\to\infty$, Lemma F.2 leads to
[…]
For $u=v$, we first observe that $\sqrt{n}(\Psi_{C,uu}-T_{uu})=\sqrt{n}(L_{C,uu}-\|\mu_{u}\|_{2}^{2})+\sqrt{n}(\sigma_{C,u}^{2}-\delta_{u}^{2}).$ For a given $1\leq j\leq p$, using the decomposition in part (b) of the proof of Theorem 3.6, we can express
$\sqrt{n}(\sigma_{C,j}^{2}-\delta_{j}^{2})=-\dfrac{2U_{j}}{\gamma_{n}}\sigma_{0j}^{2}+\mathcal{G}_{j},$
for all $1\leq j\leq p$, where $|\mathcal{G}_{j}|\lesssim(1/\sqrt{n})+(\sqrt{n}/p_{n})$ with probability at least $1-o_{P_{0}}(1).$ Letting $\mathcal{N}_{j}=-\dfrac{2U_{j}}{\gamma_{n}}\sigma_{0j}^{2},$ we have $\mathcal{N}_{j}/\sqrt{2\sigma_{0u}^{4}}\implies N(0,1)$, using a normal approximation to the Gamma distribution. Thus, we can write
$\sqrt{n}\left(\Psi_{C,uu}-T_{uu}\right)=\left(2\rho\,\sigma_{0u}\|\lambda_{0u}\|_{2}\right)R_{uu}+\mathcal{N}_{u}+\mathcal{G}_{u}.$
By construction, $R_{uu}$ is independent of $\mathcal{N}_{u}$ for $1\leq u\leq p$. Combining all these observations and using Lemmas F.1 and F.2, we have the desired result for $u=v$, with $l_{0,uu}(\rho)=2\sigma_{0u}^{4}+4\rho^{2}\sigma_{0u}^{2}\|\lambda_{0u}\|_{2}^{2}$ for $1\leq u\leq p$. ∎
Appendix D Proof of Theorem 3.8
Proof.
We first assume $u\neq v$. Let
$S_{uv}=\dfrac{n}{(n+\tau^{-2})^{2}}y^{(u)\top}UU^{\top}y^{(v)}-\lambda_{0u}^{\top}\lambda_{0v}$
and $f_{n}=n/(n+\tau^{-2})^{2}.$ Since $y^{(u)}=M_{0}\lambda_{0u}+\epsilon^{(u)}$, we can decompose $S_{uv}$ as
$S_{uv}=$ $\lambda_{0u}^{\top}(f_{n}M_{0}^{\top}UU^{\top}M_{0}-\mathbb{I}_{k})\lambda_{0v}$
$+f_{n}(\lambda_{0u}^{\top}M_{0}^{\top}UU^{\top}\epsilon^{(v)}+\lambda_{0v}^{\top}M_{0}^{\top}UU^{\top}\epsilon^{(u)}+\epsilon^{(u)\top}UU^{\top}\epsilon^{(v)}).$
Observe that $f_{n}=(1/n)+\Delta_{n}$, where $|\Delta_{n}|\asymp 1/n^{2}$. We first break up $S_{uv}$ into two parts $S_{uv}=D_{uv}+R_{uv},$ where
$D_{uv}$ $=\lambda_{0u}^{\top}\left(\dfrac{1}{n}M_{0}^{\top}U_{0}U_{0}^{\top}M_{0}-\mathbb{I}_{k}\right)\lambda_{0v}+\dfrac{1}{n}(\lambda_{0u}^{\top}M_{0}^{\top}U_{0}U_{0}^{\top}\epsilon^{(v)}+\lambda_{0v}^{\top}M_{0}^{\top}U_{0}U_{0}^{\top}\epsilon^{(u)}),$
$R_{uv}=$ $\Delta_{n}\lambda_{0u}^{\top}M_{0}^{\top}U_{0}U_{0}^{\top}M_{0}\lambda_{0v}+f_{n}\lambda_{0u}^{\top}M_{0}^{\top}(UU^{\top}-U_{0}U_{0}^{\top})M_{0}\lambda_{0v}$
$+\Delta_{n}(\lambda_{0u}^{\top}M_{0}^{\top}U_{0}U_{0}^{\top}\epsilon^{(v)}+\lambda_{0v}^{\top}M_{0}^{\top}U_{0}U_{0}^{\top}\epsilon^{(u)})$
$+f_{n}(\lambda_{0u}^{\top}M_{0}^{\top}(UU^{\top}-U_{0}U_{0}^{\top})\epsilon^{(v)}+\lambda_{0v}^{\top}M_{0}^{\top}(UU^{\top}-U_{0}U_{0}^{\top})\epsilon^{(u)})$
$+f_{n}\epsilon^{(u)\top}UU^{\top}\epsilon^{(v)}.$
We first deal with $R_{uv}$ and show that $\sqrt{n}R_{uv}=o_{P_{0}}(1)$. We use throughout that $\|M_{0}\|\asymp\sqrt{n}$ and $\|\epsilon^{(u)}\|_{2}\lesssim\sqrt{n}$ with probability at least $1-o(1),$ using Lemma E.8. First consider bounding $\sqrt{n}f_{n}\epsilon^{(u)\top}UU^{\top}\epsilon^{(v)}$ as follows:
$\left|\sqrt{n}f_{n}\epsilon^{(u)\top}UU^{\top}\epsilon^{(v)}\right|$ $\leq\sqrt{n}f_{n}\left|\epsilon^{(u)\top}U_{0}U_{0}^{\top}\epsilon^{(v)}\right|+\sqrt{n}f_{n}\left|\epsilon^{(u)\top}(UU^{\top}-U_{0}U_{0}^{\top})\,\epsilon^{(v)}\right|.$
The first term is $\sqrt{n}f_{n}\left|\epsilon^{(u)}U_{0}U_{0}^{\top}\epsilon^{(v)}\right|\,=O_{P_{0}}(\sqrt{n}f_{n})=O_{P_{0}}(1/\sqrt{n})$, since $U_{0}^{\top}\epsilon^{(u)}\sim N(0,\sigma_{0u}^{2}\mathbb{I}_{k})$ and $k$ is finite. The second term is handled by observing that
$\sqrt{n}f_{n}\left|\epsilon^{(u)\top}(UU^{\top}-U_{0}U_{0}^{\top})\epsilon^{(v)}\right|$ $\leq\sqrt{n}f_{n}\|UU^{\top}-U_{0}U_{0}^{\top}\|\|\epsilon^{(u)}\|_{2}\|\epsilon^{(v)}\|_{2}$
$\lesssim n\left(\dfrac{1}{n}+\dfrac{1}{p_{n}}\right)\sqrt{n}f_{n}$
$=o_{P_{0}}(1).$
Thus, $\sqrt{n}\left|f_{n}\epsilon^{(u)\top}UU^{\top}\epsilon^{(v)}\right|=o_{P_{0}}(1)$. We now consider the remaining terms.
  • (i)
    We first observe that with probability at least $1-o(1)$,
    $\sqrt{n}\Delta_{n}\left|\lambda_{0u}^{\top}M_{0}^{\top}U_{0}U_{0}^{\top}M_{0}\lambda_{0v}\right|$ $=\sqrt{n}\Delta_{n}\left|\lambda_{0u}M_{0}^{\top}M_{0}\lambda_{0v}\right|$
    $\leq\sqrt{n}\Delta_{n}\|M_{0}^{\top}M_{0}\|||\lambda_{0u}||_{2}||\lambda_{0v}||_{2}\lesssim\dfrac{1}{\sqrt{n}}.$
  • (ii)
    Using Proposition 3.5, we next consider
    $\sqrt{n}f_{n}\left|\lambda_{0u}^{\top}M_{0}^{\top}(UU^{\top}-U_{0}U_{0}^{\top})M_{0}\lambda_{0v}\right|$ $\lesssim\sqrt{n}f_{n}\|M_{0}^{\top}(UU^{\top}-U_{0}U_{0}^{\top})M_{0}\|$
    $\leq\sqrt{n}f_{n}\|M_{0}\|^{2}\|UU^{\top}-U_{0}U_{0}^{\top}\|$
    $=o_{P_{0}}(1).$
  • (iii)
    We next observe that with probability at least $1-o(1),$
    $\sqrt{n}\Delta_{n}\left|\lambda_{0u}^{\top}U_{0}U_{0}^{\top}M_{0}^{\top}\epsilon^{(v)}+\lambda_{0v}^{\top}U_{0}U_{0}^{\top}M_{0}^{\top}\epsilon^{(u)}\right|$ $=\sqrt{n}\Delta_{n}\left|\lambda_{0u}^{\top}M_{0}^{\top}\epsilon^{(v)}+\lambda_{0v}^{\top}M_{0}^{\top}\epsilon^{(u)}\right|$
    $\lesssim\dfrac{2}{n^{2}}(\sqrt{n})^{3}\lesssim\dfrac{1}{\sqrt{n}}.$
  • (iv)
    Using Proposition 3.5, we have lastly
    $\sqrt{n}f_{n}\left|\lambda_{0u}^{\top}M_{0}^{\top}(UU^{\top}-U_{0}U_{0}^{\top})\epsilon^{(v)}+\lambda_{0v}^{\top}M_{0}^{\top}(UU^{\top}-U_{0}U_{0}^{\top})\epsilon^{(u)}\right|$
    $\lesssim$ $\,\dfrac{2}{\sqrt{n}}\|M_{0}^{\top}\|\|UU^{\top}-U_{0}U_{0}^{\top}\|\,\|\epsilon^{(v)}\|_{2}$
    $=$ $\,o_{P_{0}}(1).$
Putting everything together implies $\sqrt{n}|R_{uv}|=o_{P_{0}}(1)$ for all $1\leq u,v\leq p$. We now look at $\sqrt{n}D_{uv}$. By the central limit theorem, we first observe that
$\sqrt{n}\lambda_{0u}^{\top}\left(\dfrac{1}{n}M_{0}^{\top}U_{0}U_{0}^{\top}M_{0}-\mathbb{I}_{k}\right)\lambda_{0v}$ $=\sqrt{n}\lambda_{0u}^{\top}\left(\dfrac{1}{n}M_{0}^{\top}M_{0}-\mathbb{I}_{k}\right)\lambda_{0v}$
$=\sqrt{n}(\overline{V}_{n}-\lambda_{0u}^{\top}\lambda_{0v})$
${\implies}N(0,\xi_{0,uv}^{2}),$
as $n\to\infty$, where $\overline{V}_{n}=(1/n)\sum_{i=1}^{n}V_{i},$ with $V_{i}:=(\lambda_{0u}^{\top}\eta_{0i})(\lambda_{0v}^{\top}\eta_{0i})$ for $i=1,\ldots,n$, and
$\begin{split}\mathcal{\xi}_{0,uv}^{2}=\mbox{var}\left[(\lambda_{0u}^{\top}\eta_{01})(\lambda_{0v}^{\top}\eta_{01})\right]&=\left\{\begin{array}[]{ll}(\lambda_{0u}^{\top}\lambda_{0v})^{2}+||\lambda_{0u}||_{2}^{2}||\lambda_{0v}||_{2}^{2},&\text{for $u\neq v$},\\ 2\|\lambda_{0u}\|_{2}^{4},&\text{for $u=v$.}\end{array}\right.\end{split}$
For $u=v$, we use the fact that $V_{i}/\|\lambda_{0u}\|_{2}^{2}$ are independent for $i=1,\ldots,n$ and marginally $\chi_{1}^{2}$ and use the central limit theorem accordingly. We next observe that for $u\neq v,$
$\sqrt{n}\left\{\dfrac{1}{n}\left(\lambda_{0u}^{\top}M_{0}^{\top}U_{0}U_{0}^{\top}\epsilon^{(v)}+\lambda_{0v}^{\top}M_{0}^{\top}U_{0}U_{0}^{\top}\epsilon^{(u)}\right)\right\}=\sqrt{n}\left\{\dfrac{1}{n}\left(\lambda_{0u}^{\top}M_{0}^{\top}\epsilon^{(v)}+\lambda_{0v}^{\top}M_{0}^{\top}\epsilon^{(u)}\right)\right\}.$
Let
$l_{uv}^{2}(M_{0})=\sigma_{0v}^{2}\lambda_{0u}^{\top}\dfrac{M_{0}^{\top}M_{0}}{n}\lambda_{0u}+\sigma_{0u}^{2}\lambda_{0v}^{\top}\dfrac{M_{0}^{\top}M_{0}}{n}\lambda_{0v}.$
Since $\epsilon^{(u)}$ and $\epsilon^{(v)}$ are independent for $u\neq v$, we have
[…]
where $Z_{uv}\sim N(0,1)$ and $Z_{uv}$ is independent of $M_{0}.$ Let
$l_{0,uv}^{2}=\sigma_{0v}^{2}||\lambda_{0u}||_{2}^{2}+\sigma_{0u}^{2}||\lambda_{0v}||_{2}^{2}.$
Then,
$\sqrt{n}D_{uv}$ $=\sqrt{n}\lambda_{0u}^{\top}\left(\dfrac{1}{n}M_{0}^{\top}U_{0}U_{0}^{\top}M_{0}-\mathbb{I}_{k}\right)\lambda_{0v}+l_{uv}(M_{0})Z_{uv}$
$=\sqrt{n}\lambda_{0u}^{\top}\left(\dfrac{1}{n}M_{0}^{\top}M_{0}-\mathbb{I}_{k}\right)\lambda_{0v}+l_{0,uv}Z_{uv}+\{l_{uv}(M_{0})-l_{0,uv}\}Z_{uv}.$
Since $\|M_{0}\|\asymp\sqrt{n}$, we have $\left|l_{uv}(M_{0})+l_{0,uv}\right|=O_{P_{0}}(1)$ and $\left|l_{uv}(M_{0})+l_{0,uv}\right|\geq l_{0,uv}\geq 2c_{1}c_{2}>0$, from Assumptions 3.2 and 3.3. Thus, the third term can be handled by observing
$\{l_{uv}(M_{0})-l_{0,uv}\}Z_{uv}$ $=\dfrac{l_{uv}^{2}(M_{0})-l_{0,uv}^{2}}{l_{uv}(M_{0})+l_{0,uv}}Z_{uv}$
$=O_{P_{0}}\left(\dfrac{1}{\sqrt{n}}\right),$
as $\left\|\dfrac{M_{0}^{\top}M_{0}}{n}-\mathbb{I}_{k}\right\|=O_{P_{0}}\left(\dfrac{1}{\sqrt{n}}\right)$ from Lemma E.1. Since $Z_{uv}$ is independent of $M_{0}$, Lemma F.1 immediately implies that
[…]
Let $\mathcal{S}_{0,uv}^{2}=l_{0,uv}^{2}+\xi_{0,uv}^{2}$. Putting all the previous results back together and invoking Slutsky’s theorem, we obtain
$\sqrt{n}S_{uv}\overset{}{\implies}N(0,\mathcal{S}_{0,uv}^{2}).$
When $u=v$, we first observe that $\delta_{u}^{2}$ is approximately independent of $\|\mu_{u}\|_{2}^{2}$ by replacing $UU^{\top}$ by $U_{0}U_{0}^{\top}$, as $U_{0}U_{0}^{\top}(\mathbb{I}_{n}-U_{0}U_{0}^{\top})=\mathbf{O}_{n\times n}$, where $\mathbf{O}_{n\times n}$ denotes the $n\times n$ matrix with all entries equal to zero. Finally, we use Proposition 3.5 as earlier to argue the remainder terms are small. Using the central limit theorem for the $\chi_{n-k}^{2}$ distribution, we obtain the desired result. ∎
Appendix E Proofs of Proposition 3.5 and Related Lemmas for Theorem 3.6
E.1 Proof of Proposition 3.5
Proof.
Suppose $X_{0}=U_{0}D_{0}V_{0}^{\top}=M_{0}\Lambda_{0}^{\top}$ is the true signal. We start with $\|X_{0}^{\top}UU^{\top}X_{0}-X_{0}^{\top}U_{0}U_{0}^{\top}X_{0}\|$ and observe that
$\|X_{0}^{\top}UU^{\top}X_{0}-X_{0}^{\top}U_{0}U_{0}^{\top}X_{0}\|$ $=\|\Lambda_{0}M_{0}^{\top}(UU^{\top}-U_{0}U_{0}^{\top})M_{0}\Lambda_{0}^{\top}\|$
$\geq s_{\min}^{2}(\Lambda_{0})\|M_{0}^{\top}(UU^{\top}-U_{0}U_{0}^{\top})M_{0}\|$
$\geq s_{\min}^{2}(\Lambda_{0})s_{\min}^{2}(M_{0})\|UU^{\top}-U_{0}U_{0}^{\top}\|,$
since $\|AB\|\geq\|A\|s_{\min}(B)$ for any matrices $A,B$, with $s_{min}(B)$ denoting the smallest singular value of $B$. Both $\Lambda_{0}$ and $M_{0}$ have exactly $k$ positive singular values since the rank is $k$. Thus, we have
$\|X_{0}^{\top}UU^{\top}X_{0}-X_{0}^{\top}U_{0}U_{0}^{\top}X_{0}\|\geq s_{\min}^{2}(\Lambda_{0})s_{\min}^{2}(M_{0})\|UU^{\top}-U_{0}U_{0}^{\top}\|.$
Since the Schatten $\infty$-norm is identical to the spectral norm, Theorem 2 in Luo et al., (2021) implies
$\|X_{0}^{\top}UU^{\top}X_{0}-X_{0}^{\top}U_{0}U_{0}^{\top}X_{0}\|=\|X_{0}^{\top}UU^{\top}X_{0}-X_{0}^{\top}X_{0}\|=\|(\mathbb{I}_{n}-UU^{\top})X_{0}\|^{2}\leq 4\|E\|^{2}.$
Thus with probability at least $1-o(1)$,
$\|UU^{\top}-U_{0}U_{0}^{\top}\|\,s_{\min}^{2}(\Lambda_{0})s_{\min}^{2}(M_{0})\leq{4\|E\|^{2}}.$
Using Lemma E.8, we have $s_{\min}(M_{0})\asymp\|M_{0}\|\ \asymp\sqrt{n}$ with probability at least $1-o(1)$, since $k=o(n)$. Under Assumption 3.2, $s_{\min}(\Lambda_{0})\asymp\|\Lambda_{0}\|\asymp\sqrt{p_{n}}$. Suppose $\sigma_{\text{sum}}^{2}=\sum_{j=1}^{p}\sigma_{0j}^{2}$ and $\sigma_{\max}^{2}=\max_{1\leq j\leq p}\sigma_{0j}^{2}=\mathcal{O}(1)$ by Assumption 3.3. Lemma E.2 implies $\|E\|\lesssim(\sigma_{\text{sum}}+\sqrt{n}\sigma_{\max})$ with probability at least $1-o(1)$. Since $\sigma_{\text{sum}}\leq\sqrt{p_{n}}\sigma_{\max}$, one has $\|E\|\lesssim\sqrt{n}+\sqrt{p_{n}}$ with probability at least $1-o(1)$, as $n\to\infty$. Thus, with probability at least $1-o(1)$,
$\|UU^{\top}-U_{0}U_{0}^{\top}\|$ $\lesssim\dfrac{(\sqrt{n}+\sqrt{p_{n}})^{2}}{np_{n}}\lesssim\dfrac{1}{n}+\dfrac{1}{p_{n}},$
which proves the desired result. ∎
E.2 Other Lemmas for Theorem 1
Lemma E.1.
Let $E_{0}\in\mathbb{R}^{n_{1}\times n_{2}}$ have i.i.d. standard Gaussian entries. Then for every $t>0$, one has
$\mathbb{P}\left(\|E_{0}\|\leq\sqrt{n_{1}}+\sqrt{n_{2}}+t\right)\geq 1-2\exp(-t^{2}/2).$
$\mathbb{E}\|E_{0}^{\top}E_{0}-n_{1}\mathbb{I}_{n_{2}}\|\leq\left(n_{1}^{1/2}+n_{2}^{1/2}\right)^{2}-n_{1}.$
Proof.
Refer to Corollary 5.35 in Vershynin, (2010). ∎
Lemma E.2.
[…]
$\mathbb{P}\left\{\|E_{0}\|\geq(1+\epsilon)(\sigma_{1}+\sigma_{2})+t\right\}\leq(n_{1}\wedge n_{2})\exp\{-t^{2}/(c_{\epsilon}^{{}^{\prime}}\sigma_{*}^{2})\}.$
Proof.
Refer to Corollary 3.11 in Bandeira and Van Handel, (2016). ∎
Lemma E.3.
For any matrix $Z\in\mathbb{R}^{n_{1}\times n_{2}}$, let its $l$th largest singular value be $s_{l}(Z)$ for $l=1,\ldots,n\wedge p$. Then for all $1\leq l\leq n\wedge p,$
[…]
Proof.
Refer to Eckart and Young, (1936). ∎
Lemma E.4.
Suppose Assumptions 3.1-3.4 hold. For all $1\leq u,v\leq p,$ we have as $n\to\infty$,
$\dfrac{1}{n}y^{(u)\top}UU^{\top}y^{(v)}\overset{P_{0}}{\to}\lambda_{0u}^{\top}\lambda_{0v}$
and
$\dfrac{1}{n}\|(\mathbb{I}_{n}-UU^{\top})y^{(u)}\|_{2}^{2}\overset{P_{0}}{\to}\sigma_{0u}^{2}.$
Proof.
We first note the identity
$\dfrac{1}{n}y^{(u)\top}UU^{\top}y^{(v)}-\lambda_{0u}^{\top}\lambda_{0v}$
$=$ $\,\lambda_{0u}^{\top}\left(\dfrac{M_{0}^{\top}M_{0}}{n}-\mathbb{I}_{k}\right)\lambda_{0v}+\dfrac{1}{n}(\lambda_{0u}^{\top}M_{0}^{\top}\epsilon^{(v)}+\lambda_{0v}^{\top}M_{0}^{\top}\epsilon^{(u)})\,$
$+\dfrac{1}{n}\epsilon^{(u)\top}U_{0}U_{0}^{\top}\epsilon^{(v)}+\dfrac{1}{n}y^{(u)\top}(UU^{\top}-U_{0}U_{0}^{\top})y^{(v)}.$
he first term is bounded by $\|\lambda_{0u}\|_{2}\|\lambda_{0v}\|_{2}\|(M_{0}^{\top}M_{0}/n)-\mathbb{I}_{k}\|\lesssim 1/\sqrt{n}$ with probability at least $1-o(1)$, using Lemma E.1. The second term is handled by first observing that $\lambda_{0u}^{\top}M_{0}^{\top}\epsilon^{(v)}\mid M_{0}\sim N\left(0,\sigma_{0v}^{2}\lambda_{0u}^{\top}M_{0}^{\top}M_{0}\lambda_{0u}\right)$ and then Lemma E.1 to argue that $|\lambda_{0u}^{\top}M_{0}^{\top}\epsilon^{(v)}|\lesssim 1/\sqrt{n}$ with probability at least $1-o(1).$ The third term is seen to be $O_{P_{0}}(1/n)$ and the fourth term is seen to be $O_{P_{0}}(n^{-1}+p_{n}^{-1})$ using Proposition 3.5 along with $\|y^{(u)}\|_{2}\lesssim\sqrt{n}$, with probability at least $1-o(1)$. Therefore, one has the desired result for all $1\leq u,v\leq p$.
To derive the second result, we first observe that $\|y^{(u)}\|_{2}^{2}\sim(\|\lambda_{0u}\|_{2}^{2}+\sigma_{0u}^{2})\,\chi_{n}^{2}$. An application of the above result along with the weak law of large numbers provides that this quantity converges in probability to $\|\lambda_{0u}\|_{2}+\sigma_{0u}^{2}-\|\lambda_{0u}\|_{2}^{2}=\sigma_{0u}^{2}$ for any $1\leq u\leq p.$
Lemma E.5.
Let $B=(b_{uv})_{1\leq u,v\leq p}$ denote the matrix of coefficients as in Section 2.2. Then, $\|B\|_{\infty}=O_{P_{0}}(1)$ under Assumptions 3.1-3.4.
Proof.
First, consider $u,v$ such that $u\neq v$. Following from Assumption 3.2, there exists $W_{1}>0$ such that $\|\lambda_{0u}\|_{2}>c_{1}$ for all $1\leq u\leq p$. Let $c_{2}=\min_{u}\sigma_{0u}^{2}>0$ from Assumption 3.3. Following from Lemma E.4, we have for sufficiently large $n$ with probability at least $1-o(1)$: (1) $\|\mu_{u}\|_{2}^{2}>c_{1}/2$ and $\mathcal{V}_{u}^{2}>c_{2}/2,$ (2) $\left|\lambda_{0u}^{\top}\lambda_{0v}\right|\leq\|\lambda_{0u}\|_{2}\|\lambda_{0v}\|_{2}\leq k\|\Lambda_{0}\|_{\infty}^{2}$. This implies that with probability at least $1-o(1)$,
$b_{uv}^{2}\leq 1+\dfrac{4k^{2}\|\Lambda_{0}\|_{\infty}^{4}}{c_{1}c_{2}}<\infty.$
The result is proved analogously for $b_{uu}$ with the bound independent of $u$. This proves the result. ∎
Lemma E.6.
Suppose Assumptions 3.1-3.4 hold. For each $j=1,\ldots,p$, we have
$\delta_{j}^{2}=\dfrac{Z_{j}}{n}+F_{j},$
where $Z_{j}/\sigma_{0j}^{2}\sim\chi_{n-k}^{2}$ and $\max_{1\leq j\leq p}|F_{j}|\lesssim\dfrac{1}{n}+\dfrac{1}{p_{n}}$ with probability at least $1-o(1)$.
Proof.
Suppose the SVD of $X_{0}=M_{0}\Lambda_{0}^{\top}=U_{0}D_{0}V_{0}^{\top}$. Let $P_{U_{0}}=U_{0}(U_{0}^{\top}U_{0})^{-1}U_{0}^{\top}=U_{0}U_{0}^{\top}$ denote the projection matrix onto the column space of $U_{0}$. We can express $\delta_{j}^{2}$ as $\delta_{j}^{2}=\dfrac{Z_{j}}{n}+F_{j},$ where we have
$Z_{j}$ $=y^{(j)\top}\left(\mathbb{I}_{n}-P_{U_{0}}\right)y^{(j)},$
$F_{j}$ $=\dfrac{\gamma_{0}\delta_{0}^{2}}{\gamma_{n}}+\dfrac{1}{n}y^{(j)\top}\left(P_{U_{0}}-\dfrac{\widehat{M}\widehat{M}^{\top}}{n+\tau^{-2}}\right)y^{(j)}-\dfrac{\gamma_{0}}{n\gamma_{n}}y^{(j)\top}\left(\mathbb{I}_{n}-\dfrac{\widehat{M}\widehat{M}^{\top}}{n+\tau^{-2}}\right)y^{(j)}.$
Letting $\mathbf{Y}=M_{0}\Lambda_{0}^{\top}+E$ where $E=[\epsilon^{(1)}\ldots\epsilon^{(p)}]$, we have $(\mathbb{I}_{n}-P_{U_{0}})\mathbf{Y}=(\mathbb{I}_{n}-P_{U_{0}})E$, implying $(\mathbb{I}_{n}-P_{U_{0}})y^{(j)}=(\mathbb{I}_{n}-P_{U_{0}})\epsilon^{(j)}\sim N_{n}(0,\sigma_{0j}^{2}(\mathbb{I}_{n}-P_{U_{0}}))$. Since $\mathbb{I}_{n}-P_{U_{0}}$ is idempotent (i.e., $(\mathbb{I}_{n}-P_{U_{0}})^{2}=\mathbb{I}_{n}-P_{U_{0}}$), we have
$\dfrac{Z_{j}}{\sigma_{0j}^{2}}=\dfrac{\|\left(\mathbb{I}_{n}-P_{U_{0}}\right)\,y^{(j)}\|_{2}^{2}}{\sigma_{0j}^{2}}\sim\chi_{\mbox{tr}(\mathbb{I}_{n}-P_{U_{0}})}^{2}\equiv\chi_{n-k}^{2}.$
We now obtain the stated probabilistic upper bound on $|F_{j}|.$ We observe
$\max_{1\leq j\leq p}|F_{j}|\leq\dfrac{\gamma_{0}\delta_{0}^{2}}{\gamma_{n}}+\dfrac{1}{n}\left\|P_{U_{0}}-\dfrac{nUU^{\top}}{n+\tau^{-2}}\right\|\max_{1\leq j\leq p}||y^{(j)}||_{2}^{2}+\dfrac{\gamma_{0}}{n\gamma_{n}}\left\|\mathbb{I}_{n}-\dfrac{nUU^{\top}}{n+\tau^{-2}}\right\|\max_{1\leq j\leq p}||y^{(j)}||_{2}^{2},$
as $\widehat{M}\widehat{M}^{\top}=nUU^{\top}$. Next, we have $\max_{1\leq j\leq p}||y^{(j)}||_{2}^{2}\lesssim n$ with probability at least $1-o(1)$, following Lemma 1 in Laurent and Massart, (2000). We start with
$G_{1}$ $:=\left\|P_{U_{0}}-\dfrac{nUU^{\top}}{n+\tau^{-2}}\right\|$
$\leq\left\|U_{0}U_{0}^{\top}-UU^{\top}\right\|+\left\|UU^{\top}-\dfrac{nUU^{\top}}{n+\tau^{-2}}\right\|$
$\lesssim\dfrac{1}{n}+\dfrac{1}{p_{n}}+\dfrac{1}{n}\lesssim\dfrac{1}{n}+\dfrac{1}{p_{n}},$
with probability at least $1-o(1)$. The second inequality is obtained using Proposition 3.5. Next, we work with
$G_{2}:=$ $\,\left\|\mathbb{I}_{n}-\dfrac{nUU^{\top}}{n+\tau^{-2}}\right\|$
$\leq$ $\,\|\mathbb{I}_{n}-UU^{\top}\|+\dfrac{\tau^{-2}}{n}\|UU^{\top}\|$
$\lesssim$  $\,1$
Combining the bounds for $G_{1}$ and $G_{2}$, we have with probability at least $1-o(1),$ $\max_{1\leq j\leq p}|F_{j}|\lesssim\dfrac{1}{n}+\dfrac{1}{p_{n}}.$
Lemma E.7.
Let $V_{n}\sim G(\gamma_{n}/2,1)$ such that $\gamma_{n}\asymp n$ and let $U_{n}=V_{n}-(\gamma_{n}/2)$. For any $a_{n}$ satisfying $a_{n}\to 0$ and $\sqrt{n}a_{n}\to\infty$,
$P(|U_{n}|\geq\gamma_{n}a_{n})\lesssim\exp(-na_{n}^{2}).$
Proof.
Immediate from Theorem 5 in Zhang and Zhou, (2020). ∎
Lemma E.8.
Let $M_{0}\in\mathbb{R}^{n\times k}$ be a matrix of iid $N(0,1)$ entries with $n>k$ and let $s_{min}(M_{0})$ and $s_{max}(M_{0})$ be the smallest and largest singular values of $M_{0}$, respectively. Then,
$\sqrt{n}-\sqrt{k-1}\lesssim s_{min}(M_{0})\leq s_{max}(M_{0})\equiv\|M_{0}\|$ $\lesssim\sqrt{n}+\sqrt{k}$
with probability at least $1-o(1).$ In particular, if $k=o(n)$, we have $\|M_{0}\|\asymp\sqrt{n}$ with probability at least $1-o(1).$
Proof.
Refer to Sections 1.1 and 1.3 in Vershynin, (2011). ∎
Appendix F Proofs of Related Lemmas for Theorems 3.7 and 3.8
Lemma F.1.
Suppose $X_{n}=Y_{n}+Z_{n},$ where $Y_{n}\overset{}{\implies}N(0,\sigma_{1}^{2})$ and $Z_{n}\sim N(0,\sigma_{2}^{2}),$ with $Y_{n}$ independent of $Z_{n}.$ Then, $X_{n}\overset{}{\implies}N(0,\sigma_{1}^{2}+\sigma_{2}^{2}).$
Proof.
It is immediate from considering the characteristic function of $X_{n}$ and taking limits. ∎
Lemma F.2.
Suppose in the setup of Theorem 3.7, we have $T_{n}=Z_{n}+Y_{n}$ for random variables $T_{n},Z_{n},Y_{n}$ such that the pseudo-posterior of $Y_{n}$ concentrates around 00 as $n,p\to\infty$ and the pseudo-posterior of $Z_{n}$ is the $N(0,1)$ density. Then as $n,p\to\infty,$
$\sup_{x}\,\left|{\Pi}_{C}(T_{n}\leq x)-\Phi(x)\right|\overset{P_{0}}{\to}0.$
Proof.
Fix $\epsilon>0$. The concentration property of $Y_{n}$ implies that ${\Pi}_{C}(|Y_{n}|>\epsilon)\overset{P_{0}}{\to}0$ as $n,p\to\infty.$ Fix $x\in\mathbb{R}$. Using the triangle inequality, one obtains
$\left|{\Pi}_{C}(T_{n}\leq x)-\Phi(x)\right|$ $\leq\left|{\Pi}_{C}(T_{n}\leq x,|Y_{n}|\leq\epsilon)-\Phi(x)\right|+{\Pi}_{C}(T_{n}\leq x,|Y_{n}|>\epsilon),$
with the second term bounded by ${\Pi}_{C}(|Y_{n}|>\epsilon)\overset{P_{0}}{\to}0.$ The first term may be decomposed as
$\left|{\Pi}_{C}(T_{n}\leq x,|Y_{n}|\leq\epsilon)-\Phi(x)\right|$ $=S_{1}+S_{2},$
where $S_{1}=|{\Pi}_{C}(T_{n}\leq x,|Y_{n}|\leq\epsilon)-\Phi(x)|\mathbbm{1}(Y_{n}>0)$ and $S_{2}=|{\Pi}_{C}(T_{n}\leq x,|Y_{n}|\leq\epsilon)-\Phi(x)|\mathbbm{1}(Y_{n}\leq 0).$ We first consider
$S_{1}=$ $\left|{\Pi}_{C}(T_{n}\leq x,|Y_{n}|\leq\epsilon)-\Phi(x)\right|\mathbbm{1}(Y_{n}>0)$
$\leq$ $\left|{\Pi}_{C}(Z_{n}\leq x-Y_{n},Y_{n}\leq\epsilon)-{\Pi}_{C}(Z_{n}\leq x,Y_{n}\leq\epsilon)\right|\mathbbm{1}(Y_{n}>0)+$
${\Pi}_{C}(Z_{n}\leq x,Y_{n}>\epsilon)\mathbbm{1}(Y_{n}>0)$
$\leq$ $\,{\Pi}_{C}(x-Y_{n}\leq Z_{n}\leq x,Y_{n}\leq\epsilon)\mathbbm{1}(Y_{n}>0)+{\Pi}_{C}(|Y_{n}|>\epsilon),$
where the inequality in the second line above is obtained by observing $\Phi(x)=\Pi_{C}(Z_{n}\leq x)=\Pi_{C}(Z_{n}\leq x,Y_{n}\leq\epsilon)+\Pi_{C}(Z_{n}\leq x,Y_{n}>\epsilon)$ and then using the triangle inequality. The first term can be bounded by $\Phi(x)-\Phi(x-\epsilon)\leq L\epsilon$ for a universal constant $L>0$, as $\Phi$ is Lipschitz continuous. The second term goes to 00 in $P_{0}$-probability, and thus is less than $L\epsilon$ with probability at least $1-o(1)$ for sufficiently large $n$. Thus, for sufficiently large $n$, $S_{1}\leq 2L\epsilon$ with probability at least $1-o(1)$. Similarly, $S_{2}\leq 2L\epsilon$ with probability at least $1-o(1)$ for sufficiently large $n$. Combining, we have
$\sup_{x}\,\left|{\Pi}_{C}(T_{n}\leq x)-\Phi(x)\right|\leq 4L\epsilon$
with probability at least $1-o(1),$ for any fixed $\epsilon>0$. This shows the result. We remark here that the result may be easily extended to the case where $Z_{n}\implies N(0,1)$ as $n\to\infty$ analogously. ∎